### v1

# Stability of time-periodic solutions to the Navier-Stokes-Fourier system

Abstract
We prove the existence and stability of a time-periodic solution to the Navier-Stokes-Fourier system in the three-dimensional whole space when a time-periodic external force is sufficiently small. The time decay estimate of the perturbation around the time-periodic solution is derived under the assumption that an initial perturbation is small and belongs to $L^{p}$ for some $1\leq p\leq 2$.
1 Introduction
The motion of a compressible viscous and heat conducting fluid is described by the Navier-Stokes-Fourier system for the density $\rho$, the velocity $v=(v_{1},v_{2},v_{3})$ and the temperature $\theta$:
$\left\{\begin{aligned} &\partial_{t}\rho+\mathrm{div}(\rho v)=0,\\ &\partial_{t}v+v\cdot\nabla v=\frac{\mu}{\rho}\Delta v+\frac{\mu+\mu^{\prime}}{\rho}\nabla\mathrm{div}\,v-\frac{\nabla P}{\rho}+f,\\ &\partial_{t}\theta+v\cdot\nabla\theta+\frac{\theta\partial_{\theta}P}{c_{V}\rho}\mathrm{div}\,v=\frac{\kappa}{c_{V}\rho}\Delta\theta+\frac{\Psi}{c_{V}\rho}.\end{aligned}\right$. (1)
Here, $f$ is the external body force, $P$ is the pressure, $\mu$ and $\mu^{\prime}$ are the viscous coefficients, $\kappa$ is the heat conductivity coefficient, $c_{V}$ is the specific heat at constant volume and $\Psi=\Psi(v)$ is the dissipation function given by
$\Psi=\frac{\mu}{2}(\partial_{i}v^{j}+\partial_{j}v^{i})^{2}+\mu^{\prime}(\partial_{i}v^{i})^{2}.$
Throughout this paper, we assume that the coefficients $\mu$, $\mu^{\prime}$, $\kappa$ and $c_{V}$ are constants which satisfy $\kappa,c_{V}>0$ and the thermodynamic restrictions
$\mu>0,\ \ \ \frac{2}{3}\mu+\mu^{\prime}\geq 0.$
The pressure $P$ is assumed to be a smooth function of $\rho$ and $\theta$ satisfying the stability conditions:
$\partial_{\rho}P>0\ \ \ \mathrm{and}\ \ \ \partial_{\theta}P>0.$
The aim of this paper is twofold. The first is to investigate the existence of time-periodic solutions to the Navier-Stokes-Fourier system in $\mathbb{R}^{3}$ subject to a given small time-periodic external force. The second is to examine the stability of the time-periodic solutions with respect to external forces and initial perturbations, respectively. Let $f$ be a time-periodic force with time-period $T>0$, that is,
$f(t,x)=f(t+T,x)\ \ \ \mathrm{for\ any}\ \ \ t\in\mathbb{R},\ x\in\mathbb{R}^{3}.$
A flow generated by $f$ is then described as the solution to the following time-periodic problem:
$\left\{\begin{aligned} &\partial_{t}\rho_{T}+\mathrm{div}(\rho_{T}v_{T})=0\ \ \ \mathrm{in}\ \ \ \mathbb{R}\times\mathbb{R}^{3},\\ &\partial_{t}v_{T}+v_{T}\cdot\nabla v_{T}=\frac{\mu}{\rho_{T}}\Delta v_{T}+\frac{\mu+\mu^{\prime}}{\rho_{T}}\nabla\mathrm{div}\,v_{T}-\frac{\nabla P}{\rho_{T}}+f\ \ \ \mathrm{in}\ \ \ \mathbb{R}\times\mathbb{R}^{3},\\ &\partial_{t}\theta_{T}+v_{T}\cdot\nabla\theta_{T}+\frac{\theta_{T}\partial_{\theta}P}{c_{V}\rho_{T}}\mathrm{div}\,v_{T}=\frac{\kappa}{c_{V}\rho_{T}}\Delta\theta_{T}+\frac{\Psi}{c_{V}\rho_{T}}\ \ \ \mathrm{in}\ \ \ \mathbb{R}\times\mathbb{R}^{3},\\ &(\rho_{T},v_{T},\theta_{T})(t,x)=(\rho_{T},v_{T},\theta_{T})(t+T,x)\ \ \ \mathrm{for}\ \ \ (t,x)\in\mathbb{R}\times\mathbb{R}^{3}.\end{aligned}\right$. (2)
We consider the time-periodic problem (2) around the constant state
$(\rho_{\infty},0,\theta_{\infty})\ \ \ \mathrm{with}\ \ \ \rho_{\infty},\theta_{\infty}>0.$
The time-periodic problem of the Navier-Stokes-Fourier system was considered by Ma, Ukai and Yang [MR2595722] in $\mathbb{R}^{N}$, $N\geq 5$. They proved the existence and stability result for the time-periodic problem by effectively applying the energy method in the inhomogeneous Sobolev space of $L^{2}$ type, but the dimensional assumption $N\geq 5$ is needed to control the $L^{2}$ norm of the time-periodic solution. Indeed, since the diffusion terms
$\mu\Delta v+(\mu+\mu^{\prime})\nabla\mathrm{div}v\ \ \ \mathrm{and}\ \ \ \frac{\kappa}{c_{V}}\Delta\theta$
are classified as elliptic operators, the spatial decay orders of the velocity $v_{T}$ and the temperature $\theta_{T}-\theta_{\infty}$ are expected at most $O(1/|x|^{N-2})$ as $|x|\to\infty$, respectively, by comparing the spatial decay order of the Newtonian potential. See also the asymptotic expansion result for the incompressible Navier-Stokes equation in [MR1794517]. If the dimension $N=3$, then the expected spatial decay order is $O(1/|x|)$, which does not belong to $L^{2}$, so that this slow spatial decay suggests a need for a new function framework to construct a time-periodic solution in dimension $N=3$. In order to establish a function space setting that admits slow spatial decay, such as $1/|x|$, we introduce the homogeneous Besov spaces as follows. Let $s\in\mathbb{R}$ and $1\leq r\leq\infty$. Choose $\phi\in C^{\infty}$ supported in the annulus $\mathcal{C}=\{\xi\in\mathbb{R}^{3}\mid 3/4\leq|\xi|\leq 8/3\}$ such that
$\sum_{j\in\mathbb{Z}}\phi(2^{-j}\xi)=1\ \ \ \mathrm{for}\ \ \ \xi\neq 0.$
We define dyadic blocks $(\dot{\Delta}_{j})_{j\in\mathbb{Z}}$ by the Fourier multiplier
$\dot{\Delta}_{j}u\equiv\mathcal{F}^{-1}[\phi(2^{-j}\cdot)\hat{u}$].
Then, the homogeneous Besov space $\dot{B}^{s}_{2,r}=\dot{B}^{s}_{2,r}(\mathbb{R}^{3})$ is defined by
$\dot{B}^{s}_{2,r}=\{u\in\mathcal{S}^{\prime}/\mathcal{P}\mid\|u\|_{\dot{B}^{s}_{2,r}}<\infty\}$
with norm
$\|u\|_{\dot{B}^{s}_{2,r}}:=\|\{2^{js}\|\dot{\Delta}_{j}u\|_{L^{2}}\}_{j\in\mathbb{Z}}\|_{\ell^{r}(\mathbb{Z})},$
where $\mathcal{S}^{\prime}$ is the set of all Schwartz distributions and $\mathcal{P}$ is the polynomial ring over $\mathbb{R}$. We shall employ a function space to construct the time-periodic solution as the space $\dot{B}^{1/2}_{2,\infty}\cap\dot{H}^{k}$, $k\in\mathbb{Z}_{\geq 5}$, which spaces allow for slow decay, such as $1/|x|$, and have good compatibility with the energy method, since the space $\dot{B}^{1/2}_{2,\infty}\cap\dot{H}^{k}$ is $L^{2}$-based space. Here, we mention that the space $\dot{B}^{1/2}_{2,1}\cap\dot{H}^{3}$ was employed for the perturbation analysis of stationary solutions to the isentropic compressible Navier-Stokes equations in [MR4803477].
Let us state our main theorem, which shows the existence and stability of the time-periodic problem (2).
Theorem 1.1.
Let $T>0$ and $k\in\mathbb{Z}_{\geq 5}$. There exists a small $\delta>0$ such that if a time-periodic force $f$ with time period $T>0$ satisfies
$\|f\|_{C(\mathbb{R};\dot{B}^{-\frac{3}{2}}_{2,\infty}\cap\dot{H}^{k-1})}\leq\delta,$
then the following assertions hold.
  1. ((i))
    (Existence of time-periodic solution) There exists a unique time-periodic solution $(\rho_{T},v_{T},\theta_{T})$ satisfying
    $\|(\rho_{T}-\rho_{\infty},v_{T},\theta_{T}-\theta_{\infty})\|_{C(\mathbb{R};\dot{B}^{\frac{1}{2}}_{2,\infty}\cap\dot{H}^{k})}\lesssim\delta.$
  2. ((ii))
    (Stability) If initial data $(\rho_{0},v_{0},\theta_{0})$ satisfy
    $\|(\rho_{0}-\rho_{\infty},v_{0},\theta_{0}-\theta_{\infty})\|_{\dot{B}^{\frac{1}{2}}_{2,\infty}\cap\dot{H}^{k}}\leq\delta,$
[…]
    $\|(\rho-\rho_{\infty},v,\theta-\theta_{\infty})\|_{C([0,\infty);\dot{B}^{\frac{1}{2}}_{2,\infty}\cap\dot{H}^{k})}\lesssim\delta.$
    In addition, for any $\epsilon>0$, if $\delta>0$ is sufficiently small and
    $(\rho_{0}-\rho_{T}(0),v_{0}-v_{T}(0),\theta_{0}-\theta_{T}(0))\in L^{p}$
    for some $1\leq p\leq 2$, then we have the time decay estimate:
    $\|(\rho-\rho_{T},v-v_{T},\theta-\theta_{T})(t)\|_{\dot{H}^{s}}$
    $\hskip 10.0pt\lesssim_{s}(1+t)^{-\frac{s}{2}-\frac{3}{2}\left(\frac{1}{p}-\frac{1}{2}\right)}\|(\rho_{0}-\rho_{T}(0),v_{0}-v_{T}(0),\theta_{0}-\theta_{T}(0))\|_{L^{p}\cap\dot{H}^{k}}$ (3)
    for any $3/2+\epsilon\leq s\leq 3/2-\epsilon$ with s/2+3/2(1/p1/2)>0s/2+3/2(1/p-1/2)>0.
The time-periodic problem for the incompressible Navier-Stokes equation has been extensively studied; we refer here to Maremonti [MR1107016], Kozono-Nakao [MR1373173], Yamazaki [MR1777114], Galdi-Silvestre [MR2603768], Kobayashi [MR2589952] and Kučera [MR2610568] among many others. In particular, Yamazaki [MR1777114] derived time-decay estimates for perturbations around a time-periodic solution to the incompressible Navier-Stokes equation in unbounded domains such as the whole space, the half-space, or exterior domains of dimension $N\geq 3$. For the isentropic compressible Navier-Stokes equation, the existence and stability of the time-periodic solution were studied by Valli [MR753158] in a 3D bounded domain, and by Kagei-Tsuda [MR3274764], Tsuda [MR3437858] in $\mathbb{R}^{N}$, $N\geq 3$. The existence and stability of the time-periodic solution to the Navier-Stokes-Fourier system was studied by Ma, Ukai and Yang [MR2595722] in $\mathbb{R}^{N}$, $N\geq 5$. We also mention that Feireisl et al. [MR2917121] show the existence of at least one weak time periodic solution of the Navier-Stokes-Fourier system in a 3D bounded domain for the specific pressure. Since the spatial decay order of the time-periodic solution is expected to be at most O(1/|x|){O}(1/|x|), the convection term $v\cdot\nabla v$ appearing in the Navier-Stokes equations is expected to exhibit the same spatial decay order, ${O}(1/|x|^{3})$, as the viscous term $\mu\Delta v$. Thus, we need to treat the convection terms carefully to drive the existence of a time-periodic solution even if the given external force is small enough. For this reason, Yamazaki and Tsuda transform the convection term $v\cdot\nabla v$ to the divergence forms $\mathrm{div}(v\otimes v)$ and $\mathrm{div}(\rho v\otimes v)$, respectively, by using the mass conservation low, and then establish a priori estimate for the solutions. In the Navier-Stokes-Fourier system case, the additional convection term $v\cdot\nabla\theta$ and the non-divergence form term $(\theta\partial_{\theta}P/c_{V}\rho)\mathrm{div}v$ appear in the equation of the temperature in (1). In Section 3 below, to establish a priori estimate in the homogeneous Besov spaces, we shall introduce the modified energy functional
$\mathcal{E}=\frac{1}{2}\left(1+\theta_{\infty}\frac{q_{2}}{p_{2}}\right)\rho v\cdot v+c_{V}\rho(\theta-\theta_{\infty})$
$\hskip 100.0pt+\frac{p_{1}}{2\rho_{\infty}}\left(1+\theta_{\infty}\left(\frac{q_{2}}{p_{2}}-\frac{q_{1}}{p_{1}}\right)\right)(\rho-\rho_{\infty})^{2},$ (4)
where $p_{1}=\partial_{\rho}P(\rho_{\infty},\theta_{\infty})$, $p_{2}=\partial_{\theta}P(\rho_{\infty},\theta_{\infty})$, $q_{1}=\partial_{\rho}^{2}P(\rho_{\infty},\theta_{\infty})$ and $q_{2}=\partial_{\rho}\partial_{\theta}P(\rho_{\infty},\theta_{\infty})$. By introducing the modified energy functional, let $\sigma=\rho-\rho_{\infty}$, $m=\rho v$ and $\eta=\theta-\theta_{\infty}$, then the equation (1) is transformed into the equation
$\left\{\begin{aligned} &\partial_{t}\sigma+\mathrm{div}\,m=0,\\ &\partial_{t}m-\frac{\mu}{\rho_{\infty}}\Delta m-\frac{\mu+\mu^{\prime}}{\rho_{\infty}}\nabla\mathrm{div}\,m+p_{1}\nabla\sigma+\frac{p_{2}}{c\rho_{\infty}}\nabla\mathcal{E}\\ &\hskip 20.0pt=-\frac{1}{\rho_{\infty}}\mathrm{div}(m\otimes m)-\frac{p_{2}}{\rho_{\infty}}\nabla(\sigma\mathcal{E})-p_{2}\nabla\left(\eta-\frac{1}{c_{V}\rho_{\infty}}\mathcal{E}\right)+\rho f+\mathcal{H}_{1},\\ &\partial_{t}\mathcal{E}-\frac{\kappa}{c_{V}\rho_{\infty}}\Delta\mathcal{E}+\frac{p_{2}\theta_{\infty}}{c_{V}\rho_{\infty}}\mathrm{div}\,m=-c_{V}\mathrm{div}(\eta m)\\ &\hskip 10.0pt-\left(1+\theta_{\infty}\frac{q_{2}}{p_{2}}\right)(p_{1}\mathrm{div}(\sigma v)+p_{2}\mathrm{div}(\eta v))+\left(1+\theta_{\infty}\frac{q_{2}}{p_{2}}\right)\rho_{\infty}f\cdot v+\mathcal{H}_{2},\end{aligned}\right$. (5)
where $\mathcal{H}_{1}$ and $\mathcal{H}_{2}$ are the higher order terms. Except for the external force terms and the higher-order terms, all nonlinear terms in (5) are written in divergence forms or potential forms. This transform is employed to establish a priori estimates for the initial value problem (1) in the proofs of Propositions 3.5 and 3.6 below. The proof of the time-decay estimate (((ii))) is inspired by the strategy employed in [MR4803477]. The key step of the proof is to derive the time-weighted estimate for the Duhamel integral:
$\sup_{0\leq t\leq\widetilde{T}_{1}}(1+t)^{\frac{s-s_{0}}{2}}\left\|\int_{0}^{t}e^{\tau A}\Phi(t-\tau)d\tau\right\|_{\dot{B}^{s}_{2,\infty}}$
$\hskip 50.0pt\lesssim\sup_{s_{1}\leq s\leq\frac{3}{2}-\epsilon}\ \sup_{0\leq t\leq\widetilde{T}_{1}}(1+t)^{\frac{s-s_{0}}{2}}\|\Phi(t)\|_{\dot{B}^{s-2}_{2,\infty}\cap\dot{B}^{\frac{3}{2}}_{2,\infty}}$,
where $\epsilon>0$, $3/2\leq s_{0}\leq 1/2$, $s_{1}=\max\{s_{0},0\}$, $\max\{s_{0},-3/2+\epsilon\}\leq s\leq 3/2-\epsilon$ and $e^{tA}$ is the linearized semigroup of (5) around the constant state which defined in (12) below. A similar analysis is found in Yamazaki [MR1777114].
The outline of this paper is as follows: In Section 2, we present some lemmas related to the homogeneous Besov space, which will be frequently used in the subsequent sections. Section 3 is devoted to the proof of Theorem 1.1. The proof of Theorem 1.1 is established by combining the local existence theorem with a priori estimates for the solutions to the initial value problem of (1). A priori estimate and the local existence result for the initial value problem are proved in Subsection 3.1 and Subsection 3.2, respectively. We show the uniform in time and the time-weighted a priori estimate for the solution in Proposition 3.1 and Proposition 3.2 below, respectively. A priori estimates are performed for the homogeneous Besov norms and the homogeneous Sobolev norms in a parallel way. The proof of the estimate in the homogeneous Besov spaces is carried out by using the modified energy functional introduced in (1). Here, the time-space norm estimate in (18) below for the linearized semigroup of (5) around the constant state $(\rho_{\infty},0,\theta_{\infty})$ plays a crucial role in deriving the homogeneous Besov norm estimates. (Cf. [MR1777114], [MR4803477].) The proof of the estimates in the homogeneous Sobolev norms is established by the energy method. The proof of the local existence is carried out by constructing the approximating sequence of the solution. The approximating sequence is constructed as the solution of the transport equation for the density and as the solution of the quasilinear parabolic equation for the velocity and the temperature, since the density causes the quasilinear effect in the equations of the velocity and the temperature in (1). By virtue of employment in this construction, we obtain the local existence result without the smallness assumption for initial data, unlike the local existence result in [MR4803477]. Finally, the proof of Theorem 1.2 is given in Subsection 3.3.
Notation.
The notation $A\lesssim_{\alpha}B$ means that there exists a constant $C$ depending on $\alpha$ such that $A\leq CB$. The notation $A\sim_{\alpha}B$ means that $A\lesssim_{\alpha}B$ and $B\lesssim_{\alpha}A$. We denote a commutator by [X,Y]=XYYX[X,Y]=XY-YX. We write $\mathcal{S}$ for the set of all Schwartz functions on $\mathbb{R}^{3}$, and we write $\mathcal{S}^{\prime}$ for the set of all tempered distributions on $\mathbb{R}^{3}$. The notations $\hat{\cdot}$ and $\mathcal{F}$ stand for the Fourier transform
$\hat{u}(\xi)=\mathcal{F}(u)(\xi)=\int_{\mathbb{R}^{3}}e^{-ix\cdot\xi}u(x)dx,$
and the notation $\mathcal{F}^{-1}$ denotes the inverse Fourier transform. We denote the $L^{2}(\mathbb{R}^{3})$ inner product by
$\langle u,v\rangle=\int_{\mathbb{R}^{3}}uvdx.$
Let $Z$ be a Banach space, $1\leq p\leq\infty$ and $I\subset\mathbb{R}$ be an interval. We denote by $L^{p}(I;Z)$ the space of all strongly measurable functions $f:I\to Z$ such that
$\|f\|_{L^{p}(I;Z)}=\|\|f(t)\|_{Z}\|_{L^{p}_{t}(I)}<\infty.$
Similarly, $C(I;Z)$ denotes the space of continuous functions $f:I\to Z$ equipped with the norm
$\|f\|_{C(I;Z)}=\sup_{t\in I}\|f(t)\|_{Z}<\infty.$
For $I=[0,T_{0})$ with $0<T_{0}\leq\infty$, we write $L^{p}_{T_{0}}(Z)$ and $C_{T_{0}}(Z)$ instead of $L^{p}(I;Z)$ and $C(I;Z)$, respectively. In particular, if $T_{0}=\infty$, we simply write $L^{p}(Z)$ and $C(Z)$ by omitting the subscript.
2 Preliminary
In this section, we present several lemmas related to the homogeneous Besov space, which are frequently used in this paper. The following basic properties for the homogeneous Besov space are proved in [MR2768550, Lemma 2.1, Proposition 2.22, Theorem 2.25 and Proposition 2.29].
Lemma 2.1.
Let $s,\tilde{s}\in\mathbb{R}$, $1\leq r,\tilde{r}\leq\infty$ and $u,v\in\mathcal{S}^{\prime}$.
  1. (i)
    (Derivative) For any $k\in\mathbb{Z}_{\geq 0}$, it follows that $\|\nabla^{k}u\|_{\dot{B}_{2,r}^{s}}\sim\|u\|_{\dot{B}_{2,r}^{s+k}}$.
  2. (ii)
    (Duality) Let $r^{\prime}$ be a conjugate exponent of $r$. Then, we have the following duality estimate:
    $\langle u,v\rangle\lesssim\|u\|_{\dot{B}^{s}_{p,r}}\|u\|_{\dot{B}^{-s}_{2,r^{\prime}}}\ \ \ \mathrm{and}\ \ \ \|u\|_{\dot{B}^{s}_{2,r}}\lesssim\sup_{\psi}\langle u,\psi\rangle,$
    where the supremum is taken over the Schwartz functions $\psi$ with
    $\|\psi\|_{\dot{B}^{-s}_{2,r^{\prime}}}\leq 1$
    and $0\notin\operatorname{supp}\mathcal{F}\psi$.
  3. (iii)
    (Interpolation) Let $s_{1}<s_{2}$ satisfy $s=(1-\theta)s_{1}+\theta s_{2}$ for some $0<\theta<1$. Then, the interpolation inequality
    $\|u\|_{\dot{B}^{s}_{2,r}}\lesssim_{\theta,s_{1},s_{2}}\|u\|^{1-\theta}_{\dot{B}^{s_{1}}_{2,\infty}}\|u\|^{\theta}_{\dot{B}^{s_{2}}_{2,\infty}}$ (6)
    holds.
  4. (iv)
    (Fatou property) Assume $s<3/2$ or $s=3/2$, $r=1$. If $\{u_{n}\}_{n\in\mathbb{Z}_{\geq 0}}$ is a bounded sequence in $\dot{B}_{2,r}^{s}\cap\dot{B}_{2,\tilde{r}}^{\tilde{s}}$, then there exists a subsequence of $\{u_{n}\}_{n}$ ((without relabeling)) and $u\in\dot{B}^{s}_{2,r}\cap\dot{B}_{2,\tilde{r}}^{\tilde{s}}$ such that
    $\lim_{n\to\infty}u_{n}=u\ \ \mathrm{in}\ \ \mathcal{S}^{\prime}\ \ \ \mathrm{and}\ \ \ \|u\|_{\dot{B}^{s}_{2,r}\cap\dot{B}_{2,\tilde{r}}^{\tilde{s}}}\lesssim\liminf_{n\to\infty}\|u_{n}\|_{\dot{B}^{s}_{2,r}\cap\dot{B}_{2,\tilde{r}}^{\tilde{s}}}$.
The following embedding property holds. (Cf. [MR2768550, Proposition 2.20, Propositon 2.39 and Theorem 2.40].)
Lemma 2.2.
For any $1\leq p\leq 2$, the space $L^{p}$ is continuously embedded in the space $\dot{B}^{-3(1/p-1/2)}_{2,\infty}$. For any $2\leq q\leq\infty$, the space $\dot{B}^{3(1/2-1/q)}_{2,1}$ is continuously embedded in the space $L^{q}$.
The following commutator estimate is the special case of [MR2768550, Lemma 2.100 and Remark 2.102].
Lemma 2.3.
Let 3/2<s<5/2-3/2<s<5/2, $1\leq r\leq\infty$. Then, we have
$\left\|\left(2^{js}\|[\dot{\Delta}_{j},v\partial_{k}]u\|_{L^{2}}\right)_{j\in\mathbb{Z}}\right\|_{\ell^{r}(\mathbb{Z})}\lesssim_{s}\|\nabla v\|_{\dot{B}^{\frac{3}{2}}_{2,1}}\|u\|_{\dot{B}^{s}_{2,r}}$
for any $1\leq k\leq 3$, where $u$ and $v$ are scalar functions.
We state the bilinear estimate in the homogeneous Besov space. The following lemma is a direct consequence of [MR2768550, Theorem 2.47 and Theorem 2.52].
Lemma 2.4.
Let $s_{1},s_{2}\in\mathbb{R}$ satisfy $s_{1},s_{2}<3/2$ and $s_{1}+s_{2}>0$. Let $1\leq r_{1},r_{2}\leq\infty$ satisfy $1/r_{1}+1/r_{2}=1/r$. Then, we have
$\|uv\|_{\dot{B}_{2,r}^{s_{1}+s_{2}-\frac{3}{2}}}\lesssim_{s_{1},s_{2}}\|u\|_{\dot{B}_{2,r_{1}}^{s_{1}}}\|v\|_{\dot{B}_{2,r_{2}}^{s_{2}}}.$
In the cases $s_{1}\leq 3/2$, $s_{2}<3/2$ with $s_{1}+s_{2}\geq 0$, we have
$\|uv\|_{\dot{B}_{2,\infty}^{s_{1}+s_{2}-\frac{3}{2}}}\lesssim\|u\|_{\dot{B}_{2,1}^{s_{1}}}\|v\|_{\dot{B}_{2,\infty}^{s_{2}}}.$
We use the following lemmas regarding the composition of functions. (Cf. [MR2768550, Theorem 2.61].)
Lemma 2.5.
Let $1\leq r\leq\infty$, $c>0$, $\Phi\in C^{\infty}(\mathbb{R})$ with $\Phi(0)=0$ and $\sigma\in\dot{B}^{s}_{2,r}\cap L^{\infty}$ with $s<2/3$, $1\leq\infty$ or $s=3/2$, $r=1$ satisfies
$\|\sigma\|_{L^{\infty}}\leq c.$
Then, we have
$\|\Phi(\sigma)\|_{\dot{B}_{2,r}^{s}}\lesssim_{c,\Phi}\|\sigma\|_{\dot{B}_{2,r}^{s}}.$
By applying Lemma 2.5 for the difference of composite functions
$\Phi(\sigma)-\Phi(\eta)=\int_{0}^{1}(\Phi^{\prime}(\eta+t(\sigma-\eta))-\Phi^{\prime}(0))(\sigma-\eta)dt+\Phi^{\prime}(0)(\sigma-\eta),$
we then obtain the following lemma.
Lemma 2.6.
Let $1\leq r\leq\infty$, $c>0$, $\Phi\in C^{\infty}(\mathbb{R})$ and $\sigma,\eta\in\dot{B}^{s}_{2,r}\cap\dot{B}^{3/2}_{2,1}$ with $3/2\leq s<3/2$, $1\leq r\leq\infty$ or $s=3/2$, $r=1$ satisfy
$\|(\sigma,\eta)\|_{L^{\infty}}\leq c.$
Then, we have
$\|\Phi(\sigma)-\Phi(\eta)\|_{\dot{B}^{s}_{2,r}}\lesssim_{c,\Phi}(1+\|(\sigma,\eta)\|_{\dot{B}^{\frac{3}{2}}_{2,1}})\|\sigma-\eta\|_{\dot{B}^{s}_{2,r}}.$ (7)
3 Existence and stability of time-periodic solution
This section is devoted to the proof of Theorem 1.1. In order to prove Theorem 1.1, we shall investigate the initial value problem of (1):
$\left\{\begin{aligned} &\partial_{t}\rho+\mathrm{div}(\rho v)=0,\\ &\partial_{t}v+v\cdot\nabla v=\frac{\mu}{\rho}\Delta v+\frac{\mu+\mu^{\prime}}{\rho}\nabla\mathrm{div}\,v-\frac{\nabla P}{\rho}+f,\\ &\partial_{t}\theta+v\cdot\nabla\theta+\frac{\theta\partial_{\theta}P}{c_{V}\rho}\mathrm{div}\,v=\frac{\kappa}{c_{V}\rho}\Delta\theta+\frac{\Psi}{c_{V}\rho},\\ &(\rho,v,\theta)|_{t=0}=(\rho_{0},v_{0},\theta_{0})\end{aligned}\right$. (8)
in $(0,T_{1})\times\mathbb{R}^{3}$ for given $T_{1}>0$.
3.1 A priori estimate
We start with a priori estimates for an initial value problem (8).
Proposition 3.1.
Let $k\in\mathbb{Z}_{\geq 3}$, $0<T_{1}<\infty$ and
$f\in C_{T_{1}}(\dot{B}^{-\frac{3}{2}}_{2,\infty}\cap\dot{H}^{k-1}).$
Suppose that the pair of functions $(\rho,v,\theta)$ with
$(\rho-\rho_{\infty},v,\theta-\theta_{\infty})\in C_{T_{1}}(\dot{B}^{\frac{1}{2}}_{2,\infty}\cap\dot{H}^{k})$ (9)
[…]
$\|(\rho-\rho_{\infty},v,\theta-\theta_{\infty})\|_{C_{T_{1}}(\dot{B}^{\frac{1}{2}}_{2,\infty}\cap\dot{H}^{k})}+\|f\|_{C_{T_{1}}(\dot{B}^{-\frac{3}{2}}_{2,\infty}\cap\dot{H}^{k-1})}\leq\delta_{1},$ (10)
then we obtain
$\|(\rho-\rho_{\infty},v,\theta-\theta_{\infty})\|_{C_{T_{1}}(\dot{B}^{\frac{1}{2}}_{2,\infty}\cap\dot{H}^{k})}$
$\hskip 30.0pt\lesssim\|(\rho_{0}-\rho_{\infty},v_{0},\theta_{0}-\theta_{\infty})\|_{\dot{B}^{\frac{1}{2}}_{2,\infty}\cap\dot{H}^{k}}+\|f\|_{C_{T_{1}}(\dot{B}^{-\frac{3}{2}}_{2,\infty}\cap\dot{H}^{k-1})}.$
Proposition 3.2.
[…]
$(\rho_{0,1}-\rho_{0,2},v_{0,1}-v_{0,2},\theta_{0,1}-\theta_{0,2})\in\dot{B}^{s_{0}}_{2,\infty}$
for some $3/2\leq s_{0}\leq 1/2$, then we have the time-weighted estimate
$\sup_{0\leq t\leq T_{1}}(1+t)^{\frac{s-s_{0}}{2}}\|(\rho_{1}-\rho_{2},v_{1}-v_{2},\theta_{1}-\theta_{2})(t)\|_{\dot{B}^{s}_{2,\infty}\cap\dot{H}^{4}}$
$\hskip 100.0pt\lesssim_{s}\|(\rho_{0,1}-\rho_{0,2},v_{0,1}-v_{0,2},\theta_{0,1}-\theta_{0,2})\|_{\dot{B}^{s_{0}}_{2,\infty}\cap\dot{H}^{4}}$
for any $3/2+\epsilon\leq s\leq 3/2-\epsilon$ with $s\geq s_{0}$.
The proofs of Proposition 3.1 and Proposition 3.2 below proceed in parallel. To derive a priori estimates in the homogeneous Besov spaces, we reformulate equation (1). Let $(\rho,v,\theta)$ be a solution of (1) with an external force $f$. Set
$\sigma=\rho-\rho_{\infty},\ \ m=\rho v\ \ \ \mathrm{and}\ \ \ \eta=\theta-\theta_{\infty}.$
We introduce the modified energy functional $\mathcal{E}=\mathcal{E}(\sigma,v,\eta)$ by
$\mathcal{E}=\frac{1}{2}\left(1+\theta_{\infty}\frac{q_{2}}{p_{2}}\right)m\cdot v+c_{V}\rho\eta+\frac{p_{1}}{2\rho_{\infty}}\left(1+\theta_{\infty}\left(\frac{q_{2}}{p_{2}}-\frac{q_{1}}{p_{1}}\right)\right)\sigma^{2},$
where $p_{1}=\partial_{\rho}P(\rho_{\infty},\theta_{\infty})$, $p_{2}=\partial_{\theta}P(\rho_{\infty},\theta_{\infty})$, $q_{1}=\partial_{\rho}^{2}P(\rho_{\infty},\theta_{\infty})$ and $q_{2}=\partial_{\rho}\partial_{\theta}P(\rho_{\infty},\theta_{\infty})$. Since we have the identities
$\frac{1}{2}\partial_{t}(\sigma^{2})=-\rho_{\infty}\sigma\mathrm{div}\,v-\sigma\mathrm{div}(\sigma v),$
$\frac{1}{2}\partial_{t}\left(m\cdot v\right)+\frac{1}{2}\mathrm{div}\left((m\cdot v)v\right)$
$\hskip 50.0pt=(\mu\Delta v+(\mu+\mu^{\prime})\nabla\mathrm{div}\,v)\cdot v-v\cdot\nabla P+f\cdot m,$
and
$\partial_{t}(\rho\eta)+\mathrm{div}(\eta m)+\frac{\theta\partial_{\theta}P}{c_{V}}\mathrm{div}v=\frac{\kappa}{c_{V}}\Delta\eta+\frac{\Psi}{c_{V}},$
the pair of functions $(\sigma,m,\mathcal{E})$ satisfies the system of equations
$\left\{\begin{aligned} &\partial_{t}\sigma+\mathrm{div}\,m=0,\\ &\partial_{t}m-\frac{1}{\rho_{\infty}}\mathcal{A}m+p_{1}\nabla\sigma+\frac{p_{2}}{c_{V}\rho_{\infty}}\nabla\mathcal{E}=G+\rho f,\\ &\partial_{t}\mathcal{E}-\frac{\kappa}{c_{V}\rho_{\infty}}\Delta\mathcal{E}+\frac{p_{2}\theta_{\infty}}{c_{V}\rho_{\infty}}\mathrm{div}\,m=H+\left(1+\theta_{\infty}\frac{q_{2}}{p_{2}}\right)f\cdot m,\end{aligned}\right$. (11)
where $\mathcal{A}=\mu\Delta m-(\mu+\mu^{\prime})\nabla\mathrm{div}$, $p_{1}=\partial_{\rho}P(\rho_{\infty},\theta_{\infty})$, $p_{2}=\partial_{\theta}P(\rho_{\infty},\theta_{\infty})$, the functions $G=G(\sigma,v,\eta)$ and $H=H_{1}+H_{2}=H_{1}(\sigma,v,\eta)+H_{2}(\sigma,v,\eta)$ are given by
$G=-\mathrm{div}\left(\frac{1}{\rho}m\otimes m\right)-p_{2}\nabla\mathcal{R}-\nabla Q_{1}+\mathcal{A}\left(\left(\frac{1}{\rho}-\frac{1}{\rho_{\infty}}\right)m\right),$
$H_{1}=-c_{V}\,\mathrm{div}\left(\eta m\right)-\left(1+\theta_{\infty}\frac{q_{2}}{p_{2}}\right)(p_{1}\mathrm{div}(\sigma v)+p_{2}\mathrm{div}(\eta v))$
$\hskip 70.0pt-\frac{1}{2}\left(1+\theta_{\infty}\frac{q_{2}}{p_{2}}\right)\mathrm{div}((m\cdot v)v)-\kappa\Delta\mathcal{R}$
$H_{2}=-\left(1+\theta_{\infty}\frac{q_{2}}{p_{2}}\right)v\cdot\nabla Q_{1}-Q_{2}\mathrm{div}\,v$
$\hskip 70.0pt-\frac{p_{1}}{\rho_{\infty}}\left(1+\theta_{\infty}\left(\frac{q_{2}}{p_{2}}-\frac{q_{1}}{p_{1}}\right)\right)\sigma\mathrm{div}(\sigma v)$
$\hskip 70.0pt+\left(1+\theta_{\infty}\frac{q_{2}}{p_{2}}\right)(\mathcal{A}v)\cdot v+\Psi.$
Here, the functions $Q_{1}=Q_{1}(\sigma,\eta)$, $Q_{2}=Q_{2}(\sigma,\eta)$ and $\mathcal{R}=\mathcal{R}(\sigma,v,\eta)$ are given by
$Q_{1}=P-P(\rho_{\infty},\theta_{\infty})-p_{1}\sigma-p_{2}\eta,$
$Q_{2}=\theta\partial_{\theta}P-\theta_{\infty}p_{2}-\theta_{\infty}q_{1}\sigma-(p_{2}+\theta_{\infty}q_{2})\eta$
and
$\mathcal{R}=\eta-\frac{1}{c_{V}\rho_{\infty}}\mathcal{E}.$
By rewriting the equations using the modified energy functional $\mathcal{E}$, the nonlinear terms $(G,H)$ in (11) are decomposed into $(G,H_{1})$, in which all terms are in divergence or potential form, and $H_{2}$, which consists only of higher-order terms.
We shall rewrite the equation (11) by using the linearized semigroup around the constant state. Let
$U=U(\sigma,v,\eta)=\left(\sqrt{p_{1}}\sigma,\ m,\ \sqrt{\frac{c_{V}}{\theta_{\infty}}}\mathcal{E}\right)^{\mathsf{T}}.$
Then, the Duhamel principle implies
$U(t)=e^{tA}U_{0}+\int_{0}^{t}e^{\tau A}\left[\begin{matrix}0\\ G+\rho f\\ \sqrt{\frac{c}{\theta_{\infty}}}H+\left(1+\theta_{\infty}\frac{q_{2}}{p_{2}}\right)\rho f\cdot v\end{matrix}\right](t-\tau)d\tau$
where $U_{0}=U|_{t=0}$ and $e^{tA}$ is a Fourier multiplier with symbol $e^{t\hat{A}(\xi)}$,
$\hat{A}(\xi)=\left[\begin{matrix}0&-i\sqrt{p_{1}}\xi^{\mathsf{T}}&0\\ -i\sqrt{p_{1}}\xi&-\frac{\mu}{\rho_{\infty}}|\xi|^{2}I_{3}+\frac{\mu+\mu^{\prime}}{\rho_{\infty}}\xi\otimes\xi&-i\frac{p_{2}}{\rho_{\infty}}\sqrt{\frac{\theta_{\infty}}{c_{V}}}\xi\\ 0&-\frac{p_{2}}{\rho_{\infty}}\sqrt{\frac{\theta_{\infty}}{c_{V}}}\xi^{\mathsf{T}}&-\frac{\kappa}{c\rho_{\infty}}|\xi|^{2}\end{matrix}\right],\ \ \ \xi\in\mathbb{R}^{3}.$ (12)
The $5\times 5$ matrix $\hat{A}(\xi)$ has eigenvalues $\lambda_{n}(\xi)$, $n=1,\ldots,4$, which satisfy that there exist constants $r_{0},\beta>0$, such that
$\operatorname{Re}\lambda_{n}(\xi)\sim-|\xi|^{2}\ \ \ \mathrm{for}\ \ \ |\xi|\leq r_{0}$ (13)
and
$\operatorname{Re}\lambda_{n}(\xi)\leq-\beta\ \ \ \mathrm{for}\ \ \ |\xi|>r_{0}$ (14)
for $n=1,\ldots,4$. Furthermore, the spectral resolutions
$e^{t\hat{A}(\xi)}=\sum_{n=1}^{4}e^{t\lambda_{n}(\xi)}P_{n}(\xi),\ \ \ \xi\in\mathbb{R}^{3},$ (15)
[…]
$\dot{S}_{j}u=\sum_{j^{\prime}<j}\dot{\Delta}_{j^{\prime}}u,\ \ \ u\in\mathcal{S}^{\prime}.$
[…]
Lemma 3.3.
Let $s,s_{0}\in\mathbb{R}$ satisfy $s_{0}\leq s$. Then, there exists $j_{0}\in\mathbb{Z}$ such that
$\|\dot{S}_{j_{0}}e^{tA}V_{0}\|_{\dot{B}^{s}_{2,r}}\lesssim(1+t)^{-\frac{s-s_{0}}{2}}\|V_{0}\|_{\dot{B}^{s_{0}}_{2,r}},\ \ \ t\geq 0,$ (16)
for any $1\leq r\leq\infty$, $V_{0}\in\dot{B}^{s_{0}}_{2,\infty}$ and
$\|(1-\dot{S}_{j_{0}})e^{tA}V_{0}\|_{\dot{B}^{s}_{2,r}}\lesssim e^{-ct}\|V_{0}\|_{\dot{B}^{s_{0}}_{2,r}},\ \ \ t\geq 0,$ (17)
for any $1\leq r\leq\infty$, $V_{0}\in\dot{B}^{s}_{2,\infty}$, where $c>0$ is a constant. Furthermore, for any $1\leq r\leq\infty$, we have
$\|\dot{S}_{j_{0}}e^{tA}V_{0}\|_{L^{r}(\dot{B}^{s}_{2,r})}\lesssim\|V_{0}\|_{\dot{B}^{s-\frac{2}{r}}_{2,r}},\ \ \ V_{0}\in\dot{B}^{s-\frac{2}{r}}_{2,r}.$ (18)
[…]
Proof.
From (13) and (14), it follows that there exist constants $c_{1},c_{2}>0$ and $j_{0}\in\mathbb{Z}$ such that
$|\phi(2^{-j}\xi)e^{t\lambda_{n}(\xi)}|\lesssim\begin{cases}e^{-c_{1}t|\xi|^{2}}&\text{if $j<j_{0}$,}\\ e^{-c_{2}t}&\text{if $j\geq j_{0}$.}\end{cases}$
for any $n=1,\ldots,4$ and $\xi\in\mathbb{R}^{3}$. Then, the spectral resolutions (15) yield
$2^{sj}\|\dot{\Delta}_{j}\dot{S}_{j_{0}}e^{tA}V_{0}\|_{L^{2}}\lesssim_{r_{0}}2^{sj}e^{-\tilde{c}_{1}2^{2j}t}\|\dot{\Delta}_{j}V_{0}\|_{L^{2}}$ (19)
and
$2^{sj}\|\dot{\Delta}_{j}(1-\dot{S}_{j_{0}})e^{tA}V_{0}\|_{L^{2}}\lesssim_{r_{0}}e^{-\tilde{c}_{2}t}2^{sj}\|\dot{\Delta}_{j}V_{0}\|_{L^{2}},$ (20)
where $\tilde{c}_{1},\tilde{c}_{2}>0$ are constants. By taking the $\ell^{r}(\mathbb{Z})$ norm and the $L^{r}(\ell^{r}(\mathbb{Z}))$ norm, respectively, of both sides of (19), we obtain the estimates (16) and (18), respectively. The estimate (17) is obtained by taking the $\ell^{r}(\mathbb{Z})$ norm of both sides of (20). ∎
The following lemma will be used to derive the time-weighted estimate. The proof of Lemma 3.4 below is the same as in [MR4803477, Proposition 5.2].
Lemma 3.4.
Let $\widetilde{T}_{1}>0$, $3/2\leq s_{0}\leq 1/2$ and $\Phi\in L^{\infty}_{T_{2}}(\dot{B}^{s_{1}-2}_{2,\infty}\cap\dot{B}^{3/2}_{2,\infty})$ with $s_{1}=\max\{s_{0},0\}$. For any $0<\epsilon<1/2$, we set
$D_{\epsilon}=\sup_{s_{1}\leq s\leq\frac{3}{2}-\epsilon}\ \sup_{0\leq t\leq\widetilde{T}_{1}}(1+t)^{\frac{s-s_{0}}{2}}\|\Phi(t)\|_{\dot{B}^{s-2}_{2,\infty}\cap\dot{B}^{\frac{3}{2}}_{2,\infty}}$,
then we have
$\sup_{0\leq t\leq\widetilde{T}_{1}}(1+t)^{\frac{s-s_{0}}{2}}\left\|\int_{0}^{t}e^{\tau A}\Phi(t-\tau)d\tau\right\|_{\dot{B}^{s}_{2,\infty}}\lesssim_{\epsilon}D_{\epsilon}$ (21)
for any $s\in\mathbb{R}$ with $\max\{s_{0},-3/2+\epsilon\}\leq s\leq 3/2-\epsilon$.
Proof.
We first show the estimate (21) with $s=3/2-\epsilon$. Let
$\alpha_{0}=\begin{cases}3&\text{if $-3/2\leq s_{0}\leq 0$,}\\ 5/2&\text{if $0<s_{0}\leq 1/2$.}\end{cases}$
For any test function $\psi\in\mathcal{S}$, by Lemma 2.1 (ii),
$\left\langle\dot{S}_{j_{0}}\int_{0}^{t}e^{\tau A}\Phi(t-\tau)d\tau,\psi\right\rangle$ $=\int_{0}^{t}\langle\Phi(t-\tau),\dot{S}_{j_{0}}e^{\tau A^{*}}\psi\rangle d\tau$
$\lesssim\int_{0}^{\frac{t}{2}}\|\Phi(t-\tau)\|_{\dot{B}^{-\frac{1}{2}-\epsilon}_{2,\infty}}\|\dot{S}_{j_{0}}e^{\tau A^{*}}\psi\|_{\dot{B}^{\frac{1}{2}+\epsilon}_{2,1}}d\tau$
$\hskip 5.0pt+\int_{\frac{t}{2}}^{t}\|\Phi(t-\tau)\|_{\dot{B}^{\frac{3}{2}-\epsilon-\alpha_{0}}_{2,\infty}}\|\dot{S}_{j_{0}}e^{\tau A^{*}}\psi\|_{\dot{B}^{\alpha_{0}-\frac{3}{2}+\epsilon}_{2,1}}d\tau.$
By Lemma 3.3, we have
$\|\dot{S}_{j_{0}}e^{\tau A^{*}}\psi\|_{\dot{B}^{\alpha_{0}-\frac{3}{2}+\epsilon}_{2,1}}$ $\lesssim(1+\tau)^{-\frac{\alpha_{0}}{2}}\|\psi\|_{\dot{B}^{-\frac{3}{2}+\epsilon}_{2,1}}$
$\lesssim(1+t)^{-\frac{\alpha_{0}}{2}}\|\psi\|_{\dot{B}^{-\frac{3}{2}+\epsilon}_{2,1}},\ \ \ \frac{t}{2}\leq\tau\leq t,$
and
$\int_{0}^{\frac{t}{2}}\|\dot{S}_{j_{0}}e^{\tau A^{*}}\psi\|_{\dot{B}^{\frac{1}{2}+\epsilon}_{2,1}}d\tau\lesssim\|\psi\|_{\dot{B}^{-\frac{3}{2}+\epsilon}_{2,1}},$
so that we obtain
$\left\langle\dot{S}_{j_{0}}\int_{0}^{t}e^{\tau A}\Phi(t-\tau)d\tau,\psi\right\rangle\lesssim(1+t)^{-\frac{3}{4}+\frac{s_{0}+\epsilon}{2}}D_{\epsilon}.$
Lemma 2.1 (ii) now shows that
$\left\|\dot{S}_{j_{0}}\int_{0}^{t}e^{\tau A}\Phi(t-\tau)d\tau\right\|_{\dot{B}^{\frac{3}{2}-\epsilon}_{2,\infty}}\lesssim(1+t)^{-\frac{3}{4}+\frac{s_{0}+\epsilon}{2}}D_{\epsilon}.$
By Lemma 3.3,
$\left\|(1-\dot{S}_{j_{0}})\int_{0}^{t}e^{\tau A}\Phi(t-\tau)d\tau\right\|_{\dot{B}^{\frac{3}{2}-\epsilon}_{2,\infty}}$ $\lesssim\int_{0}^{t}e^{-ct}\|\Phi(t-\tau)\|_{\dot{B}^{\frac{3}{2}-\epsilon}_{2,\infty}}d\tau$
$\lesssim(1+t)^{-\frac{3}{4}+\frac{s_{0}+\epsilon}{2}}D_{\epsilon}.$
Then, we obtain the estimate (21) with $s=3/2-\epsilon$. Next, we show the estimate (21) with $s=\max\{s_{0},-3/2+\epsilon\}$. If $s<1/2$, then by Lemma 3.3, we have
$\left\|\int_{0}^{t}e^{\tau A}\Phi(t-\tau)d\tau\right\|_{\dot{B}^{s}_{2,\infty}}$ $\lesssim\int_{0}^{t}(1+\tau)^{-\frac{3}{4}-\frac{s_{0}}{2}}\|\Phi(t-\tau)\|_{\dot{B}^{-\frac{3}{2}}_{2,\infty}\cap\dot{B}^{s}_{2,\infty}}d\tau$
$\lesssim D_{\epsilon}\int_{0}^{t}(1+\tau)^{-\frac{3}{4}-\frac{s}{2}}(1+t-\tau)^{-\frac{1}{4}+\frac{s}{2}}d\tau$
$\lesssim D_{\epsilon}.$
If $s=1/2$, then Lemma 3.3 implies that
$\left\|\int_{0}^{t}e^{\tau A}\Phi(t-\tau)d\tau\right\|_{\dot{B}^{\frac{1}{2}}_{2,\infty}}\lesssim D_{\epsilon}.$
By interpolating the estimates in (21) with $s=\max\{s_{0},-3/2+\epsilon\}$ and $s=3/2-\epsilon$, we obtain the estimate (21) for any $\max\{s_{0},-3/2+\epsilon\}\leq s\leq 3/2-\epsilon$. ∎
The following proposition shows the time-weighted estimate in $\dot{B}^{s}_{2,\infty}$ for the differences
$\tilde{\sigma}:=\rho_{1}-\rho_{2},\ \tilde{v}:=v_{1}-v_{2},\ \tilde{\eta}:=\eta_{1}-\eta_{2}.$
Proposition 3.5.
Under the assumption of Proposition 3.2 with small $\delta_{1}>0$, for any $3/2+\epsilon\leq s\leq 3/2-\epsilon$ and $3/2\leq s_{0}\leq 1/2$ with $s_{0}\leq s$, we have
$\sup_{0\leq t\leq T_{1}}(1+t)^{\frac{s-s_{0}}{2}}\|(\tilde{\sigma},\tilde{v},\tilde{\eta})(t)\|_{\dot{B}^{s}_{2,\infty}}\lesssim_{\epsilon}\delta_{1}\mathcal{D}_{\epsilon}(T_{1})+\|(\tilde{\sigma}_{0},\tilde{v}_{0},\tilde{\eta}_{0})\|_{\dot{B}^{s_{0}}_{2,\infty}\cap\dot{B}^{\frac{3}{2}-\epsilon}_{2,\infty}}$,
where $\tilde{\sigma}_{0}=\rho_{0,1}-\rho_{0,2}$, $\tilde{v}_{0}=v_{0,1}-v_{0,2}$, $\tilde{\eta}_{0}=\theta_{0,1}-\theta_{0,2}$ and
$\mathcal{D}_{\epsilon}(T_{1})=\sup_{s_{1}\leq\tilde{s}\leq\frac{3}{2}-\epsilon}\sup_{0\leq t\leq T_{1}}(1+t)^{\frac{\tilde{s}-s_{0}}{2}}\|(\tilde{\sigma},\tilde{v},\tilde{\eta})\|_{\dot{B}^{\tilde{s}}_{2,\infty}\cap\dot{H}^{4}},\ \ s_{1}=\max\{0,s_{0}\}.$
Proof.
By Lemma 2.4 and the assumption (10), if $\delta_{1}>0$ is small enough, then
$\|(\tilde{\sigma},\tilde{v},\tilde{\eta})\|_{\dot{B}^{s}_{2,\infty}}\sim_{\delta_{1},\epsilon}\|\tilde{U}\|_{\dot{B}^{s}_{2,\infty}}$
for any $3/2+\epsilon\leq s\leq 3/2-\epsilon$, where $\tilde{U}:=U(\sigma_{1},v_{1},\eta_{1})-U(\sigma_{2},v_{2},\eta_{2})$. Thus, by Lemma 3.4, it suffices to show that
$\|(\tilde{G},\tilde{H}_{1})\|_{\dot{B}^{s-2}_{2,\infty}\cap\dot{B}^{\frac{3}{2}}_{2,\infty}}\lesssim\delta_{1}\|(\tilde{\sigma},\tilde{v},\tilde{\eta})\|_{\dot{B}^{s}_{2,\infty}\cap\dot{H}^{4}}$ (22)
for any $s_{1}\leq s\leq 3/2-\epsilon$, where $\tilde{G}=G(\sigma_{1},v_{1},\eta_{1})-G(\sigma_{2},v_{2},\eta_{2})$, $\tilde{H}_{1}=H_{1}(\sigma_{1},v_{1},\eta_{1})-H_{1}(\sigma_{2},v_{2},\eta_{2})$, and
$\left\|\int_{0}^{t}e^{\tau A}\tilde{\Gamma}(t-\tau)d\tau\right\|_{\dot{B}^{s}_{2,\infty}}\lesssim\delta_{1}(1+t)^{-\frac{s-s_{0}}{2}}\mathcal{D}_{\epsilon}(T_{1})$ (23)
for any $s_{1}\leq s\leq 3/2-\epsilon$, where $\tilde{\Gamma}=\Gamma(\sigma_{1},v_{1},\eta_{1})-\Gamma(\sigma_{2},v_{2},\eta_{2})$ with
$\Gamma(\sigma,v,\eta):=\left[\begin{matrix}0\\ \rho f\\ \sqrt{\frac{c}{\theta_{\infty}}}H_{2}(\sigma,v,\eta)+\left(1+\theta_{\infty}\frac{q_{2}}{p_{2}}\right)\rho f\cdot v\end{matrix}\right].$
We first show the estimate (23). Applying Taylor’s theorem for $Q_{1}=Q_{1}(\sigma,\eta)$, $Q_{2}=Q_{2}(\sigma,\eta)$, respectively, there are smooth functions $R_{1}=R_{1}(\sigma,\eta)$, $R_{2}=R_{2}(\sigma,\eta)$ with $R_{1}(0,0)=R_{2}(0,0)=0$ such that
$Q_{1}=\frac{1}{2}[\sigma,\eta](D^{2}P(\rho_{\infty},\theta_{\infty})+R_{1})[\sigma,\eta]^{\mathsf{T}}$
and
$Q_{2}=\frac{1}{2}[\sigma,\eta](D^{2}\partial_{\theta}P(\rho_{\infty},\theta_{\infty})+R_{2})[\sigma,\eta]^{\mathsf{T}},$
where $D^{2}P$ and $D^{2}\partial_{\theta}P$ are the Hesse matrices. Since terms $v\cdot\nabla Q_{1}(\sigma,\eta)$ and $Q_{2}(\sigma,\eta)\mathrm{div}\,v$ consist of product of at least three functions, Lemma 2.4 shows that
$\|v_{1}\cdot\nabla Q_{1}(\sigma_{1},\eta_{1})-v_{2}\cdot\nabla Q_{1}(\sigma_{2},\eta_{2})\|_{\dot{B}^{-\frac{3}{2}}_{2,\infty}\cap\dot{B}^{\frac{3}{2}}_{2,\infty}}\lesssim_{\delta_{1}}\delta_{1}^{2}\|(\tilde{\sigma},\tilde{\eta})\|_{\dot{B}^{\frac{3}{2}-\epsilon}_{2,\infty}\cap\dot{H}^{3}}$
and
$\|Q_{2}(\sigma_{1},\eta_{1})\mathrm{div}\,v_{1}-Q_{2}(\sigma_{2},\eta_{2})\mathrm{div}\,v_{2}\|_{\dot{B}^{-\frac{3}{2}}_{2,\infty}\cap\dot{B}^{\frac{3}{2}}_{2,\infty}}$ $\lesssim_{\delta_{1}}\delta_{1}^{2}\|(\tilde{\sigma},\tilde{v},\tilde{\eta})\|_{\dot{B}^{\frac{3}{2}-\epsilon}_{2,\infty}\cap\dot{H}^{3}}.$
Thus, Lemma 2.4 shows
$\|H_{2}(\sigma_{1},v_{1},\eta_{1})-H_{2}(\sigma_{2},v_{2},\eta_{2})\|_{\dot{B}^{-\frac{3}{2}}_{2,\infty}\cap\dot{B}^{\frac{3}{2}}_{2,\infty}}\lesssim\delta_{1}\|(\tilde{\sigma},\tilde{v},\tilde{\eta})\|_{\dot{B}^{\frac{3}{2}-\epsilon}_{2,\infty}\cap\dot{H}^{3}}.$
By Lemma 2.4 and Lemma 3.3, we have
$\left\|\int_{0}^{t}e^{tA}\tilde{\Gamma}(t-\tau)d\tau\right\|_{\dot{B}^{s}_{2,\infty}}\lesssim\int^{t}_{0}(1+\tau)^{-\frac{3}{4}-\frac{s}{2}}\|\tilde{\Gamma}(t-\tau)\|_{\dot{B}^{-\frac{2}{3}}_{2,\infty}\cap\dot{B}^{s}_{2,\infty}}d\tau$
$\hskip 50.0pt\lesssim\int_{0}^{t}(1+\tau)^{-\frac{3}{4}-\frac{s}{2}}\|\tilde{\Gamma}(t-\tau)\|_{\dot{B}^{-\frac{2}{3}}_{2,\infty}\cap\dot{B}^{s}_{2,\infty}}d\tau$
$\hskip 50.0pt\lesssim(\|f\|_{\dot{B}^{-\frac{3}{2}+\epsilon}_{2,1}}+\delta_{1}^{2})\int_{0}^{t}(1+\tau)^{-\frac{3}{4}-\frac{s}{2}}\|(\tilde{\sigma},\tilde{v},\tilde{\eta})\|_{\dot{B}^{\frac{3}{2}-\epsilon}_{2,\infty}\cap\dot{H}^{3}}d\tau$
$\hskip 50.0pt\lesssim\delta_{1}\mathcal{D}_{\epsilon}(T_{1})(1+t)^{-\frac{s-s_{0}}{2}}.$
We proceed to show that the estimate in (22). By Lemma 2.4, for any $s_{1}\leq s\leq 3/2-\epsilon$, we have
$\left\|\mathrm{div}\left(\rho_{1}v_{1}\otimes v_{1}\right)-\mathrm{div}\left(\rho_{2}v_{2}\otimes v_{2}\right)\right\|_{\dot{B}^{s-2}_{2,\infty}\cap\dot{B}^{\frac{3}{2}}_{2,\infty}}$
$\hskip 30.0pt\lesssim\left\|\rho_{1}v_{1}\otimes v_{1}-\rho_{2}v_{2}\otimes v_{2}\right\|_{\dot{B}^{s-1}_{2,\infty}\cap\dot{B}^{\frac{5}{2}}_{2,\infty}}$
$\hskip 30.0pt\lesssim\delta_{1}\|(\tilde{\sigma},\tilde{v})\|_{\dot{B}^{s}_{2,\infty}\cap\dot{H}^{3}},$
$\left\|\nabla\mathcal{R}(\sigma_{1},v_{1},\eta_{1})-\nabla\mathcal{R}(\sigma_{2},v_{2},\eta_{2})\right\|_{\dot{B}^{s-2}_{2,\infty}\cap\dot{B}^{\frac{3}{2}}_{2,\infty}}\lesssim\delta_{1}\|(\tilde{\sigma},\tilde{v},\tilde{\eta})\|_{\dot{B}^{s}_{2,\infty}\cap\dot{H}^{3}},$
$\left\|\Delta\mathcal{R}(\sigma_{1},v_{1},\eta_{1})-\Delta\mathcal{R}(\sigma_{2},v_{2},\eta_{2})\right\|_{\dot{B}^{s-2}_{2,\infty}\cap\dot{B}^{\frac{3}{2}}_{2,\infty}}$
$\hskip 30.0pt\lesssim\left\|\mathcal{R}(\sigma_{1},v_{1},\eta_{1})-\mathcal{R}(\sigma_{2},v_{2},\eta_{2})\right\|_{\dot{B}^{s}_{2,\infty}\cap\dot{B}^{\frac{7}{2}}_{2,\infty}}$
$\hskip 30.0pt\lesssim\delta_{1}\|(\tilde{\sigma},\tilde{v},\tilde{\eta})\|_{\dot{B}^{s}_{2,\infty}\cap\dot{H}^{4}},$
and
$\|\mathrm{div}((\rho_{1}v_{1}\cdot v_{1})v_{1})-\mathrm{div}((\rho_{2}v_{2}\cdot v_{2})v_{2})\|_{\dot{B}^{s-2}_{2,\infty}\cap\dot{B}^{\frac{3}{2}}_{2,\infty}}$
$\hskip 30.0pt\lesssim\|(\rho_{1}v_{1}\cdot v_{1})v_{1}-(\rho_{2}v_{2}\cdot v_{2})v_{2}\|_{\dot{B}^{s-1}_{2,\infty}\cap\dot{B}^{\frac{5}{2}}_{2,\infty}}$
$\hskip 30.0pt\lesssim\delta_{1}\|(\tilde{\sigma},\tilde{v})\|_{\dot{B}^{s}_{2,\infty}\cap\dot{H}^{3}}$
The estimates for the remaining terms in $\tilde{G}$ and $\tilde{H}_{1}$ follow similarly. ∎
By repeating the proof of Proposition 3.5 for $(\sigma,v,\eta)$ in the case $s=s_{0}=1/2$, we obtain the following proposition.
Proposition 3.6.
Under the assumption of Proposition 3.1 with small $\delta_{1}>0$, we have
$\|(\sigma,v,\eta)\|_{C_{T_{1}}(\dot{B}^{\frac{1}{2}}_{2,\infty})}\lesssim\delta_{1}\|(\sigma,v,\eta)\|_{C_{T_{1}}(\dot{B}^{\frac{1}{2}}_{2,\infty}\cap\dot{H}^{3})}$
$\hskip 60.0pt+\|(\sigma_{0},v_{0},\eta_{0})\|_{\dot{B}^{\frac{1}{2}}_{2,\infty}}+\|f\|_{C_{T_{1}}(\dot{B}^{-\frac{3}{2}}_{2,\infty}\cap\dot{B}^{-\frac{1}{2}}_{2,\infty})},$
where $\sigma_{0}=\rho_{0}-\rho_{\infty}$ and $\eta_{0}=\theta_{0}-\theta_{\infty}$.
The estimates in the homogeneous Sobolev spaces $\dot{H}^{k}$ are shown using the energy method for a solution to the initial value problem (8).
Proposition 3.7.
Under the assumption of Proposition 3.1 with small $\delta_{1}>0$, we have
$\|(\sigma,v,\eta)\|_{C_{T_{1}}(\dot{H}^{1}\cap\dot{H}^{k})}\lesssim\|(\sigma,v,\eta)\|_{C_{T_{1}}(\dot{B}^{\frac{1}{2}}_{2,\infty})}$
$\hskip 60.0pt+\|(\sigma_{0},v_{0},\eta_{0})\|_{\dot{H}^{1}\cap\dot{H}^{k}}+\|f\|_{C_{T_{1}}({H}^{k-1})}.$ (24)
Proof.
Since $(\sigma,v,\eta)$ satisfies (1), we have the energy identities:
$\frac{1}{2}\partial_{t}\left\|\nabla^{\ell}\left(\frac{p_{1}}{\rho_{\infty}^{2}}\sigma,\,v,\,\frac{c}{\theta_{\infty}}\eta\right)\right\|_{L^{2}}^{2}$
$+\frac{\mu}{\rho_{\infty}}\|\nabla^{\ell+1}v\|_{L^{2}}^{2}+\frac{\mu+\mu^{\prime}}{\rho_{\infty}}\|\nabla^{\ell}\mathrm{div}\,v\|_{L^{2}}^{2}+\frac{\kappa}{\rho_{\infty}\theta_{\infty}}\|\nabla^{\ell+1}\eta\|_{L^{2}}^{2}$
$=\frac{p_{1}}{\rho_{\infty}^{2}}\langle\nabla^{\ell}\mathrm{div}(\sigma v),\nabla^{\ell}\sigma\rangle+\langle\nabla^{\ell}R_{1},\nabla^{\ell}v\rangle+\left\langle\frac{1}{\theta}\nabla^{\ell}R_{2},\nabla^{\ell}\eta\right\rangle$ (25)
for $\ell=1,\ldots,k$, and
$\partial_{t}\langle\nabla^{\ell-1}v,\nabla^{\ell-1}\nabla\sigma\rangle+\frac{P^{\prime}(\rho_{\infty})}{\rho_{\infty}}\|\nabla^{\ell}\sigma\|_{L^{2}}^{2}$
$=\rho_{\infty}\|\nabla^{\ell-1}\mathrm{div}\,v\|_{L^{2}}^{2}-\langle\nabla^{\ell-1}v,\nabla^{\ell-1}\nabla\mathrm{div}(\sigma v)\rangle$
$\hskip 70.0pt+\left\langle\nabla^{\ell-1}\left(\frac{1}{\rho_{\infty}}\mathcal{A}v+R_{1}\right),\nabla^{\ell-1}\nabla\sigma\right\rangle,$ (26)
for any $\ell=2,\ldots,k$, where the functions $R_{1}=R_{1}(\sigma,v,\eta)$, $R_{2}=R_{2}(\sigma,v,\eta)$ and $R_{3}=R_{3}(\sigma,v,\eta)$ are given by
$R_{1}=-v\cdot\nabla v+\left(\frac{1}{\rho}-\frac{1}{\rho_{\infty}}\right)\mathcal{A}v$
$\hskip 50.0pt-\left(\frac{1}{\rho}-\frac{1}{\rho_{\infty}}\right)(p_{1}\nabla\sigma+p_{2}\nabla\eta)-\frac{\nabla Q_{1}}{\rho}+f$
and
$R_{2}=-cv\cdot\nabla\eta-\left(\frac{\theta\partial_{\theta}P}{\rho}-\frac{\theta_{\infty}p_{2}}{\rho_{\infty}}\right)\mathrm{div}\,v+\kappa\left(\frac{1}{\rho}-\frac{1}{\rho_{\infty}}\right)\Delta\theta+\frac{\Psi}{\rho}.$
For any multi-index $\alpha\in\mathbb{Z}_{\geq 0}^{3}$ with $|\alpha|=k$,
$\langle\partial^{\alpha}_{x}\mathrm{div}(\sigma v),\partial_{x}^{\alpha}\sigma\rangle=-\frac{1}{2}\langle\mathrm{div}\,v\partial^{\alpha}_{x}\sigma,\partial^{\alpha}_{x}\sigma\rangle+\sum_{0\leq\beta<\alpha}\langle\mathrm{div}(\partial^{\beta}_{x}\sigma\partial^{\alpha-\beta}_{x}v),\partial^{\alpha}_{x}\sigma\rangle,$ (27)
so that we obtain
$|\langle\nabla^{\ell}\mathrm{div}(\sigma v),\nabla^{\ell}\sigma\rangle|$
$\lesssim\|v\|_{\dot{H}^{1}\cap\dot{H}^{k+1}}\|\sigma\|_{\dot{H}^{2}\cap\dot{H}^{k}}^{2}+\|v\|_{\dot{H}^{2}\cap\dot{H}^{k+1}}\|\sigma\|_{\dot{H}^{1}\cap\dot{H}^{k}}\|\sigma\|_{\dot{H}^{2}\cap\dot{H}^{k}},$ (28)
for any $\ell=1,\ldots,k$. Sobolev’s inequality implies that
$\|R_{1}\|_{\dot{H}^{\ell-1}}+\|R_{2}\|_{\dot{H}^{\ell-1}}$
$\hskip 20.0pt\lesssim\|(v,\eta)\|_{\dot{H}^{2}\cap\dot{H}^{k+1}}\|(\sigma,v,\eta)\|_{\dot{H}^{1}\cap\dot{H}^{k+1}}+\|f\|_{H^{k-1}}$ (29)
for $\ell=1,\ldots,k$. For $\gamma>0$, we define the energy function
$E_{k}(t):=\sum_{\ell=1}^{k}\left\|\nabla^{\ell}\left(\frac{p_{1}}{\rho_{\infty}^{2}}\sigma,\,v,\,\frac{c}{\theta_{\infty}}\eta\right)(t)\right\|_{L^{2}}^{2}+\gamma\sum_{\ell=2}^{k}\langle\nabla^{\ell-1}v(t),\nabla^{\ell-1}\nabla\sigma(t)\rangle$
for $0\leq t\leq T_{1}$. By Lemma 2.4, we have
$E_{k}\lesssim\|(\sigma,v,\eta)\|_{\dot{B}^{\frac{1}{2}}_{2,\infty}}^{2}+\|(v,\eta)\|_{\dot{H}^{2}\cap\dot{H}^{k+1}}^{2}+\|\sigma\|_{\dot{H}^{2}\cap\dot{H}^{k}}^{2}.$
for $0\leq t\leq T_{1}$. Combining (3.1), (3.1), (3.1) and (3.1), if $\delta_{1}>0$ and $\gamma>0$ are small enough, then there is a constant $c>0$ such that
$\frac{d}{dt}E_{k}+cE_{k}\lesssim\|(\sigma,v,\eta)\|_{C_{T}(\dot{B}^{\frac{1}{2}}_{2,\infty})}^{2}+\|f\|_{C_{T_{1}}(H^{k-1})}^{2}\ \ \ \mathrm{for}\ \ \ 0\leq t\leq T_{1}.$
Therefore, Grönwall’s inequality implies
$E_{k}(t)$ $\lesssim e^{-ct}E_{k}(0)+\int_{0}^{t}e^{-c\tau}d\tau(\|(\sigma,v,\eta)\|_{C_{T_{1}}(\dot{B}^{\frac{1}{2}}_{2,\infty})}^{2}+\|f\|_{C_{T_{1}}(H^{k-1})}^{2})$
$\lesssim E_{k}(0)+\|(\sigma,v,\eta)\|_{C_{T_{1}}(\dot{B}^{\frac{1}{2}}_{2,\infty})}^{2}+\|f\|_{C_{T_{1}}(H^{k-1})}^{2}\ \ \ \mathrm{for}\ \ \ 0\leq t\leq T_{1},$
so that we obtain the estimate (3.7) when $\gamma>0$ is small. ∎
Propositions 3.6 and 3.7 now lead to Proposition 3.1. The following proposition shows the time-weighted estimate for the $\dot{H}^{4}$ norm. The proof is carried out in a similar manner to that of Proposition 3.7.
Proposition 3.8.
Under the assumption of Proposition 3.2, we have the estimate
$\sup_{0\leq t\leq T_{1}}(1+t)^{\frac{3}{4}-\frac{\epsilon+s_{0}}{2}}\|(\tilde{\sigma},\tilde{v},\tilde{\eta})(t)\|_{\dot{H}^{4}}$
$\hskip 60.0pt\lesssim_{\epsilon}\tilde{D}(T_{1})+\delta_{1}\mathcal{D}_{\epsilon}(T_{1})+\|(\tilde{\sigma}_{0},\tilde{v}_{0},\tilde{\eta}_{0})\|_{\dot{B}^{s_{0}}_{2,\infty}\cap\dot{H}^{4}},$ (30)
where
$\tilde{D}(T_{1})=\sup_{0\leq t\leq T_{1}}(1+t)^{\frac{3}{4}-\frac{\epsilon+s_{0}}{2}}\|(\tilde{\sigma},\tilde{v},\tilde{\eta})(t)\|_{\dot{B}^{\frac{3}{2}-\epsilon}_{2,\infty}}$
Proof.
Since $(\sigma_{1},v_{1},\eta_{1})$ and $(\sigma_{2},v_{2},\eta_{2})$ satisfy (1), respectively, we have the energy estimate for $(\tilde{\sigma},\tilde{v},\tilde{\eta})=(\sigma_{1},v_{1},\eta_{1})-(\sigma_{2},v_{2},\eta_{2})$:
$\frac{1}{2}\partial_{t}\left\|\nabla^{4}\left(\frac{p_{1}}{\rho_{\infty}^{2}}\tilde{\sigma},\,\tilde{v},\,\frac{c}{\theta_{\infty}}\tilde{\eta}\right)\right\|_{L^{2}}^{2}$
$+\frac{\mu}{\rho_{\infty}}\|\nabla^{5}\tilde{v}\|_{L^{2}}^{2}+\frac{\mu+\mu^{\prime}}{\rho_{\infty}}\|\nabla^{4}\mathrm{div}\,\tilde{v}\|_{L^{2}}^{2}+\frac{\kappa}{\rho_{\infty}\theta_{\infty}}\|\nabla^{5}\tilde{\eta}\|_{L^{2}}^{2}$
$=\frac{p_{1}}{\rho_{\infty}^{2}}\langle\nabla^{4}(\mathrm{div}(\tilde{\sigma}v_{1}+\sigma_{2}\tilde{v})),\nabla^{4}\tilde{\sigma}\rangle+\langle\nabla^{4}\tilde{R}_{1},\nabla^{4}\tilde{v}\rangle$
$\hskip 30.0pt+\left\langle\nabla^{4}\left(\left(\frac{1}{\theta_{1}}-\frac{1}{\theta_{2}}\right){R}_{2}\right),\nabla^{4}\tilde{\eta}\right\rangle+\left\langle\nabla^{4}\left(\frac{1}{\theta_{2}}\tilde{R}_{2}\right),\nabla^{4}\tilde{\eta}\right\rangle$, (31)
and
$\partial_{t}\langle\nabla^{3}\tilde{v},\nabla^{3}\nabla\tilde{\sigma}\rangle+\frac{P^{\prime}(\rho_{\infty})}{\rho_{\infty}}\|\nabla^{4}\tilde{\sigma}\|_{L^{2}}^{2}$
[…]
$\hskip 70.0pt+\left\langle\nabla^{3}\left(\frac{1}{\rho_{\infty}}\mathcal{A}\tilde{v}+\tilde{R_{1}}\right),\nabla^{3}\nabla\tilde{\sigma}\right\rangle$, (32)
where $\tilde{R}_{1}=R_{1}(\sigma_{1},v_{1},\eta_{1})-R_{1}(\sigma_{2},v_{2},\eta_{2})$, $\tilde{R}_{2}=R_{2}(\sigma_{1},v_{1},\eta_{1})-R_{2}(\sigma_{2},v_{2},\eta_{2})$. By Sobolev’s inequality and Lemma 2.4, we have
$\langle\nabla^{4}\tilde{R}_{1},\nabla^{4}\tilde{v}\rangle+\left\langle\nabla^{4}\left(\left(\frac{1}{\theta_{1}}-\frac{1}{\theta_{2}}\right){R}_{2}\right),\nabla^{4}\tilde{\eta}\right\rangle+\left\langle\nabla^{4}\left(\frac{1}{\theta_{2}}\tilde{R}_{2}\right),\nabla^{4}\tilde{\eta}\right\rangle$
$\lesssim\delta_{1}\|(\tilde{\sigma},\tilde{v},\tilde{\eta})\|_{\dot{B}^{\frac{3}{2}-\epsilon}_{2,\infty}\cap\dot{H}^{4}}\|\nabla^{5}(\tilde{v},\tilde{\eta})\|_{L^{2}}+\delta_{1}\|\nabla^{5}(\tilde{v},\tilde{\eta})\|_{L^{2}}^{2}$
and
$\langle\nabla^{3}\tilde{v},\nabla^{3}\nabla\mathrm{div}(\tilde{\sigma}v_{1}+\sigma_{2}\tilde{v})\rangle+\left\langle\nabla^{3}\left(\frac{1}{\rho_{\infty}}\mathcal{A}\tilde{v}+\tilde{R_{1}}\right),\nabla^{3}\nabla\tilde{\sigma}\right\rangle$
$\lesssim\delta_{1}\|(\tilde{\sigma},\tilde{v})\|_{\dot{B}^{\frac{3}{2}-\epsilon}_{2,\infty}\cap\dot{H}^{3}}\|(\nabla^{5}\tilde{v},\nabla^{4}\tilde{\sigma})\|_{L^{2}}+\|\nabla^{5}\tilde{v}\|_{L^{2}}\|\nabla^{4}\tilde{\sigma}\|_{L^{2}}.$
By using the identity (27), we have
$\langle\nabla^{4}(\mathrm{div}(\tilde{\sigma}v_{1}+\sigma_{2}\tilde{v})),\nabla^{4}\tilde{\sigma}\rangle$
$\hskip 30.0pt\lesssim\delta_{1}(\|(\tilde{\sigma},\tilde{v})\|_{\dot{B}^{\frac{3}{2}-\epsilon}_{2,\infty}\cap\dot{H}^{4}}+\|\nabla^{5}\tilde{v}\|_{L^{2}})\|\nabla^{4}\tilde{\sigma}\|_{L^{2}}.$
Therefore, there exists a constant $c>0$ such that the energy
$E_{\gamma}(t):=\left\|\nabla^{4}\left(\frac{p_{1}}{\rho_{\infty}^{2}}\tilde{\sigma},\,\tilde{v},\,\frac{c}{\theta_{\infty}}\tilde{\eta}\right)(t)\right\|_{L^{2}}^{2}+\gamma\langle\nabla^{3}\tilde{v}(t),\nabla^{3}\nabla\tilde{\sigma}(t)\rangle,\ \ \ 0\leq t\leq T_{1}$
satisfies the differential inequality
$\frac{d}{dt}E_{\gamma}+cE_{\gamma}\lesssim(\tilde{D}(T_{1})+\delta_{1}\mathcal{D}_{\epsilon}(T_{1}))^{2}(1+t)^{-\frac{3}{2}+{\epsilon+s_{0}}}$
[…]
Now, Proposition 3.2 is obtained by Proposition 3.5 and Proposition 3.8.
3.2 Local existence
We will use the following local existence result to prove Theorem 1.1. The proof is carried out by establishing the local existence and estimates for both the transport equation and the quasilinear parabolic equation, and then combining these to construct a sequence of approximate solutions converging to the local solution.
Proposition 3.9.
Let $k\in\mathbb{Z}_{\geq 3}$ Then there exist a constant $c_{2}$ with $0<c_{2}<1$ such that if $(\sigma_{0},v_{0},\eta_{0})\in\dot{B}^{1/2}_{2,\infty}\cap\dot{H}^{k}$,
$\|\sigma_{0}\|_{\dot{B}^{\frac{3}{2}}_{2,1}}\leq c_{2}\rho_{\infty}$
[…]
$\|(\sigma,v,\eta)\|_{C_{T_{2}}(\dot{B}^{\frac{1}{2}}_{2,\infty}\cap\dot{H}^{k})}\leq 2\|(\sigma_{0},v_{0},\eta_{0})\|_{\dot{B}^{\frac{1}{2}}_{2,\infty}\cap\dot{H}^{k}}$
and
$\|\sigma\|_{C_{T_{2}}(L^{\infty})}\leq\tilde{c}_{2}\rho_{\infty}.$
Proof.
Lemma 2.2, there is a constant $C_{2}>1$ such that
$\|\tilde{\sigma}\|_{L^{\infty}}\leq C_{2}\|\tilde{\sigma}\|_{\dot{B}^{\frac{3}{2}}_{2,1}}\ \ \ \mathrm{for\ any}\ \ \ \tilde{\sigma}\in\dot{B}^{\frac{3}{2}}_{2,1}.$
Let $c_{2}=1/(2C_{2})$ and assume that an initial value $\sigma_{0}$ satisfies
$\|\sigma_{0}\|_{\dot{B}^{\frac{3}{2}}_{2,1}}\leq c_{2}\rho_{\infty}.$
Then, we have
$\inf_{x\in\mathbb{R}^{3}}\rho_{0}(x)\geq\frac{1}{2}\rho_{\infty}.$
For fixed $T_{2}$ with $0<T_{2}\leq 1$, which we will choose later, we construct the approximate sequence $\{(\rho_{n},v_{n},\eta_{n})\}_{n\geq 0}=\{(\rho_{\infty}+\sigma_{n},v_{n},\eta_{n})\}_{n\geq 0}$ with
$\{(\sigma_{n},v_{n},\eta_{n})\}_{n\geq 0}\subset C_{T_{2}}(\dot{B}^{1/2}_{2,\infty}\cap\dot{H}^{k})$
by induction. Let $n\in\mathbb{Z}_{>0}$. Assume $(\sigma_{n-1},v_{n-1},\eta_{n-1})\in C_{T_{2}}(\dot{B}^{1/2}_{2,\infty}\cap\dot{H}^{k})$ satisfies $(v_{n-1},\eta_{n-1})\in C_{T_{2}}(\dot{H}^{k+1})$,
$\|(\sigma_{n-1},v_{n-1},\eta_{n-1})\|_{C_{T_{2}}(\dot{B}^{\frac{1}{2}}_{2,\infty}\cap\dot{H}^{k})}+\|(v_{n-1},\eta_{n-1})\|_{L^{2}_{T_{2}}(\dot{H}^{k+1})}$
$\hskip 70.0pt\leq 2\|(\sigma_{0},v_{0},\eta_{0})\|_{\dot{B}^{\frac{1}{2}}_{2,\infty}\cap\dot{H}^{k}}=:2E_{0}$
and
$\|\sigma_{n-1}\|_{C_{T_{2}}(\dot{B}^{\frac{3}{2}}_{2,\infty})}\leq\tilde{c}_{2}\rho_{\infty},$
where $\tilde{c}_{2}:=(1+c_{2})/2$. Let $\sigma_{n}$ be a solution of the initial value problem of the transport equation
$\left\{\begin{aligned} &\partial_{t}\sigma_{n}+v_{n-1}\cdot\nabla\sigma_{n}=\tilde{F}_{n-1}\ \ \ \mathrm{in}\ \ \ (0,T_{2})\times\mathbb{R}^{3},\\ &\sigma_{n}|_{t=0}=\sigma_{0}\ \ \ \mathrm{on}\ \ \ \mathbb{R}^{3},\end{aligned}\right$. (33)
where
$\tilde{F}_{n-1}=-\sigma_{n-1}\mathrm{div}\,v_{n-1}.$
By the identities
$-\langle v_{n-1}\cdot\nabla\nabla^{k}\sigma_{n},\nabla^{k}\sigma_{n}\rangle=\frac{1}{2}\langle\mathrm{div}\,v_{n-1}\nabla^{k}\sigma_{n},\nabla^{k}\sigma_{n}\rangle$
and, for any $j\in\mathbb{Z}$,
$-2^{j}\langle\dot{\Delta}_{j}(v_{n-1}\cdot\nabla\sigma_{n}),\dot{\Delta}_{j}\sigma_{n}\rangle$
$\hskip 50.0pt=\frac{1}{2}2^{j}\langle\mathrm{div}\,v_{n-1}\dot{\Delta}_{j}\sigma_{n},\dot{\Delta}_{j}\sigma_{n}\rangle-2^{j}\langle[\dot{\Delta}_{j},v_{n-1}\cdot\nabla]\sigma_{n},\dot{\Delta}_{j}\sigma_{n}\rangle,$
it follows from Lemma 2.4 that, for any $0\leq t\leq T_{2}$ and $\alpha\in\mathbb{Z}^{3}_{\geq 0}$,
$\|\partial^{\alpha}_{x}\sigma_{n}(t)\|_{L^{2}}^{2}-\|\partial^{\alpha}_{x}\sigma_{0}\|_{L^{2}}^{2}=\int_{0}^{t}\langle\mathrm{div}\,v_{n-1}\partial^{\alpha}_{x}\sigma_{n}+2\partial^{\alpha}_{x}\tilde{F}_{n-1},\partial^{\alpha}_{x}\sigma_{n}\rangle d\tau$
$\hskip 100.0pt+\sum_{0<\beta\leq\alpha}\int_{0}^{t}\langle\partial_{x}^{\beta}v_{n-1}\cdot\partial_{x}^{\alpha-\beta}\nabla\sigma_{n},\partial_{x}^{\alpha}\sigma_{n}\rangle d\tau$
$\leq T_{2}\|\mathrm{div}\,v_{n-1}\|_{C_{T_{2}}(\dot{H}^{1}\cap\dot{H}^{k})}\|\sigma_{n}\|^{2}_{C_{T_{2}}(\dot{H}^{1}\cap\dot{H}^{k})}+T_{2}E_{0}^{2}\|\partial^{\alpha}_{x}\sigma_{n}\|_{C_{T_{2}}(L^{2})},$
$\|\sigma_{n}(t)\|_{\dot{B}^{\frac{3}{2}}_{2,1}}^{2}-\|\sigma_{0}\|_{\dot{B}^{\frac{3}{2}}_{2,1}}^{2}=\sum_{j\in\mathbb{Z}}2^{3j}\int_{0}^{t}\langle\dot{\Delta}_{j}(v_{n-1}\cdot\nabla\sigma_{n}+2\tilde{F}_{n-1}),\dot{\Delta}_{j}\sigma_{n}\rangle d\tau$
$\hskip 5.0pt\lesssim\|\nabla v_{n-1}\|_{L^{1}_{T_{2}}(\dot{B}^{\frac{3}{2}}_{2,1})}\|\sigma_{n}\|_{C_{T_{2}}(\dot{B}^{\frac{3}{2}}_{2,1})}^{2}$
$\hskip 20.0pt+(T_{2}E_{0}+T_{2}^{\frac{1}{2}}\|v_{n-1}\|_{L^{2}_{T_{2}}(\dot{H}^{k+1})})\|\sigma_{n-1}\|_{C_{T_{2}}(\dot{B}^{\frac{3}{2}}_{2,1})}\|\sigma_{n}\|_{C_{T_{2}}(\dot{B}^{\frac{3}{2}}_{2,1})}$
$\hskip 5.0pt\lesssim T_{2}^{\frac{1}{2}}E_{0}\|\sigma_{n}\|_{C_{T_{2}}(\dot{B}^{\frac{3}{2}}_{2,1})}^{2}+T_{2}^{\frac{1}{2}}E_{0}\|\sigma_{n-1}\|_{C_{T_{2}}(\dot{B}^{\frac{3}{2}}_{2,1})}\|\sigma_{n}\|_{C_{T_{2}}(\dot{B}^{\frac{3}{2}}_{2,1})}$
and
$\|\sigma_{n}\|_{C_{T_{2}}(\dot{B}^{\frac{1}{2}}_{2,\infty})}^{2}-\|\sigma_{0}\|_{\dot{B}^{\frac{1}{2}}_{2,\infty}}^{2}\lesssim T_{2}^{\frac{1}{2}}E_{0}\|\sigma_{n}\|_{C_{T_{2}}(\dot{B}^{\frac{1}{2}}_{2,\infty})}^{2}+T_{2}E_{0}^{2}\|\sigma_{n}\|_{C_{T_{2}}(\dot{B}^{\frac{1}{2}}_{2,\infty})},$
so that there exist constants $\tilde{c}>0$ such that if
$T_{2}(E_{0}+E_{0}^{2})\leq\tilde{c}$,
then we obtain
$\|\sigma_{n}\|_{C_{T_{2}}(\dot{B}^{\frac{1}{2}}_{2,\infty}\cap\dot{H}^{k})}\leq 2E_{0}$
and
$\|\sigma_{n}\|_{C_{T_{2}}(\dot{B}^{\frac{3}{2}}_{2,1})}\leq\tilde{c_{2}}\rho_{\infty}.$ (34)
Let $(v_{n},\eta_{n})$ be a solution of the initial value problem
$\left\{\begin{aligned} &\partial_{t}v_{n}-\frac{1}{\rho_{n-1}}\mathcal{A}v_{n}=\tilde{G}_{n-1}\ \ \ \mathrm{in}\ \ \ (0,T_{2})\times\mathbb{R}^{3},\\ &\partial_{t}\eta_{n}-\frac{\kappa}{c\rho_{\infty}}\Delta\eta_{n}=\tilde{H}_{n-1}\ \ \ \mathrm{in}\ \ \ (0,T_{2})\times\mathbb{R}^{3},\\ &(v_{n},\eta_{n})|_{t=0}=(v_{0},\eta_{0})\ \ \ \mathrm{on}\ \ \ \mathbb{R}^{3}\end{aligned}\right$. (35)
where
$\tilde{G}_{n-1}=-v_{n-1}\cdot\nabla v_{n-1}-\frac{\nabla(P(\rho_{\infty}+\sigma_{n-1},\theta_{\infty}+\eta_{n-1}))}{\rho_{\infty}+\sigma_{n-1}}+f$
and
$\tilde{H}_{n-1}=-v_{n-1}\cdot\nabla\eta_{n-1}$
$\hskip 20.0pt-\frac{(\theta_{\infty}+\eta_{n-1})\partial_{\theta}P(\rho_{\infty}+\sigma_{n-1},\theta_{\infty}+\eta_{n-1})}{c(\rho_{\infty}+\sigma_{n-1})}\mathrm{div}v_{n-1}+\frac{\Psi(v_{n-1})}{1+\sigma_{n-1}}.$
Then, for any $0\leq t\leq T_{2}$,
$\|\nabla^{k}v_{n}(t)\|_{L^{2}}^{2}-\|\nabla^{k}v_{0}\|_{L^{2}}^{2}$
$\hskip 50.0pt+\int_{0}^{t}\frac{\mu}{\rho_{\infty}+\sigma_{n-1}}\|\nabla^{k+1}v_{n}\|_{L^{2}}^{2}+\frac{\mu+\mu^{\prime}}{\rho_{\infty}+\sigma_{n-1}}\|\nabla^{k+1}v_{n}\|_{L^{2}}^{2}d\tau$
$\hskip 10.0pt\lesssim\int_{0}^{t}(\|\sigma_{n-1}\|_{\dot{H}^{1}\cap\dot{H}^{k}}\|v_{n}\|_{\dot{H}^{1}\cap\dot{H}^{k}}+\|\tilde{G}_{n-1}\|_{\dot{H}^{k-1}})\|\nabla^{k+1}v_{n}\|_{L^{2}}d\tau,$
$\|\nabla^{k}\eta_{n}(t)\|_{L^{2}}^{2}-\|\nabla^{k}\eta_{0}\|_{L^{2}}^{2}+\int_{0}^{t}\frac{\kappa}{c(\rho_{\infty}+\sigma_{n-1})}\|\nabla^{k+1}\eta_{n}\|_{L^{2}}^{2}d\tau$
$\hskip 10.0pt\lesssim\int_{0}^{t}(\|\sigma_{n-1}\|_{\dot{H}^{1}\cap\dot{H}^{k}}\|\eta_{n}\|_{\dot{H}^{1}\cap\dot{H}^{k}}+\|\tilde{H}_{n-1}\|_{\dot{H}^{k-1}})\|\nabla^{k+1}\eta_{n}\|_{L^{2}}d\tau.$
For any $j\in\mathbb{Z}$ and $t$ with $0\leq t\leq T_{2}$,
$2^{j}\|\dot{\Delta}_{j}(v_{n},\eta_{n})(t)\|_{L^{2}}^{2}-2^{j}\|\dot{\Delta}_{j}(v_{0},\eta_{0})\|_{L^{2}}^{2}$
$\hskip 30.0pt+2^{j}\int_{0}^{t}\mu\|\nabla\dot{\Delta}_{j}v_{n}\|_{L^{2}}^{2}+(\mu+\mu^{\prime})\|\mathrm{div}\dot{\Delta}_{j}v_{n}\|_{L^{2}}^{2}+\frac{k}{c}\|\nabla\dot{\Delta}_{j}\eta_{n}\|_{L^{2}}^{2}d\tau$
$\hskip 10.0pt\lesssim\int_{0}^{t}\|\sigma_{n-1}\|_{\dot{B}^{\frac{1}{2}}_{2,\infty}}\|(v_{n},\eta_{n})\|_{\dot{B}^{\frac{1}{2}}_{2,\infty}\cap\dot{H}^{k+1}}2^{\frac{1}{2}j}\|\nabla\dot{\Delta}_{j}(v_{n},\eta_{n})\|_{L^{2}}d\tau$
$\hskip 130.0pt+\int_{0}^{t}\|(\tilde{G}_{n-1},\tilde{H}_{n-1})\|_{L^{2}}2^{j}\|\dot{\Delta}_{j}(v_{n},\eta_{n})\|_{L^{2}}d\tau.$
Since
$\|(\tilde{G}_{n-1},\tilde{H}_{n-1})\|_{{H}^{k-1}}\lesssim E_{0}+E_{0}^{2},$
there is a constant $C>0$ such that
$\|(v_{n},\eta_{n})\|_{C_{T_{2}}(\dot{B}^{\frac{1}{2}}_{2,\infty}\cap\dot{H}^{k})}+\|(v_{n},\eta_{n})\|_{L^{2}_{T_{2}}(\dot{H}^{k+1})}$
$\hskip 50.0pt\leq\|(\sigma_{0},v_{0},\eta_{0})\|_{\dot{B}^{\frac{1}{2}}_{2,\infty}\cap\dot{H}^{k}}+T_{0}^{\frac{1}{4}}(E_{0}+E_{0}^{2}).$
Therefore, if we take $\tilde{c}=\tilde{c}(E_{0})>0$ small enough, then we have
$\|(\sigma_{n},v_{n},\eta_{n})\|_{C_{T_{2}}(\dot{B}^{\frac{1}{2}}_{2,\infty}\cap\dot{H}^{k})}+\|(v_{n},\eta_{n})\|_{L^{2}_{T_{2}}(\dot{H}^{k+1})}\leq 2E_{0}$ (36)
when $T_{2}^{\frac{1}{4}}(E_{0}+E_{0}^{2})\leq\tilde{c}$. Now, we obtain the sequence $\{(\rho_{n},v_{n},\eta_{n})\}_{n\geq 0}=\{(\rho_{\infty}+\sigma_{n},v_{n},\eta_{n})\}_{n\geq 0}$
$\{(\sigma_{n},v_{n},\eta_{n})\}_{n\geq 0}\subset C_{T_{2}}(\dot{B}^{1/2}_{2,\infty}\cap\dot{H}^{k})$
for fixed $T_{2}>0$ with $T_{2}^{\frac{1}{4}}(E_{0}+E_{0}^{2})\leq\tilde{c}$, which satisfies (33), (34), (35) and (36) for any $n\in\mathbb{Z}_{>0}$. Let $\delta\sigma_{n}=\sigma_{n}-\sigma_{n-1}$, $\delta v_{n}=v_{n}-v_{n-1}$ and $\delta\eta_{n}=\eta_{n}-\eta_{n-1}$ for $n\in\mathbb{Z}_{>0}$. We next show the contractivity estimate
$\|(\delta\sigma_{n},\delta v_{n},\delta\eta_{n})\|_{C_{T_{2}}(\dot{B}^{\frac{1}{2}}_{2,\infty})}\leq\frac{1}{2}\|(\delta\sigma_{n-1},\delta v_{n-1},\delta\eta_{n-1})\|_{C_{T_{2}}(\dot{B}^{\frac{1}{2}}_{2,\infty})},\ \ n\in\mathbb{Z}_{\geq 2}$ (37)
for small $T_{2}>0$. The pair of functions $(\delta v_{n},\delta\eta_{n})$ is a solution of the initial value problem
$\left\{\begin{aligned} &\partial_{t}\delta v_{n}-\frac{1}{\rho_{n-1}}\mathcal{A}\delta v_{n}\\ &\hskip 30.0pt=-\frac{\delta\sigma_{n-1}}{\rho_{n-1}\rho_{n-2}}\mathcal{A}v_{n-1}+\delta\tilde{G}_{n-1}\ \ \ \mathrm{in}\ \ \ (0,T_{2})\times\mathbb{R}^{3},\\ &\partial_{t}\delta\eta_{n}-\frac{\kappa}{c\rho_{n-1}}\Delta\delta\eta_{n}=\delta\tilde{H}_{n-1}\ \ \ \mathrm{in}\ \ \ (0,T_{2})\times\mathbb{R}^{3},\\ &(\delta v_{n},\delta\eta_{n})|_{t=0}=(0,0)\ \ \ \mathrm{on}\ \ \ \mathbb{R}^{3},\end{aligned}\right$. (38)
where $\delta G_{n-1}=\tilde{G}_{n-1}-\tilde{G}_{n-2}$, $\delta\tilde{H}_{n-1}=\tilde{H}_{n-1}-\tilde{H}_{n-2}$. By Lemma 2.3, for any $j\in\mathbb{Z}$,
$2^{-\frac{1}{2}j}\left\|\left[\dot{\Delta}_{j},\rho_{n-1}^{-1}\mathrm{div}\right]\nabla\delta v_{n}+\left[\dot{\Delta}_{j},\rho_{n-1}^{-1}\nabla\right]\mathrm{div}\delta v_{n}\right\|_{L^{2}}$
$\lesssim\|\nabla\sigma_{n-1}\|_{\dot{B}^{\frac{3}{2}}_{2,1}}\|\delta v_{n}\|_{\dot{B}^{\frac{1}{2}}_{2,\infty}}.$
By (34), for any $j\in\mathbb{Z}$,
$2^{j}\left\langle\frac{1}{\rho_{n-1}}\mathcal{A}\dot{\Delta}_{j}\delta v_{n},\dot{\Delta}_{j}\delta v_{n}\right\rangle$
$\hskip 40.0pt\gtrsim 2^{j}\|\nabla\dot{\Delta}_{j}\delta v_{n}\|_{L^{2}}^{2}-\|\nabla\sigma_{n-1}\|_{L^{\infty}}^{2}\|\delta v_{n}\|_{\dot{B}^{\frac{1}{2}}_{2,\infty}}^{2}.$
By Lemma 2.4, we have
$\left\|\frac{\delta\sigma_{n-1}}{\rho_{n-1}\rho_{n-2}}\mathcal{A}v_{n-1}\right\|_{\dot{B}^{-\frac{1}{2}}_{2,\infty}}\lesssim\|\delta\sigma_{n-1}\|_{\dot{B}^{\frac{1}{2}}_{2,\infty}}\|v_{n}\|_{\dot{B}^{\frac{5}{2}}_{2,\infty}}.$
Then, for any $j\in\mathbb{Z}$ and $0\leq t\leq T_{2}$, we have
$2^{j}\|\dot{\Delta}_{j}\delta v_{n}(t)\|_{L^{2}}^{2}+\int_{0}^{t}2^{j}\|\nabla\dot{\Delta}_{j}\delta v_{n}\|_{L^{2}}^{2}d\tau$
$\lesssim_{E_{0}}\int_{0}^{t}\|\nabla\sigma_{n-1}\|_{\dot{B}^{\frac{3}{2}}_{2,1}}^{2}\|\delta v_{n}\|_{\dot{B}^{\frac{1}{2}}_{2,\infty}}^{2}+\left\|\frac{\delta\sigma_{n-1}}{\rho_{n-1}\rho_{n-2}}\mathcal{A}v_{n-1}\right\|_{\dot{B}^{-\frac{1}{2}}_{2,\infty}}^{2}$
$\hskip 150.0pt+\|\delta\tilde{G}_{n-1}\|_{\dot{B}^{-\frac{1}{2}}_{2,\infty}}^{2}d\tau$
$\lesssim T_{2}E_{0}^{2}\|(\delta v_{n},\delta\sigma_{n-1})\|_{C_{T_{2}}(\dot{B}^{\frac{1}{2}}_{2,\infty})}^{2}+T_{2}\|\delta\tilde{G}_{n-1}\|_{C_{T_{2}}(\dot{B}^{-\frac{1}{2}}_{2,\infty})}^{2}.$
Since $\delta\eta_{n}$ is a solution of (38), for any $j\in\mathbb{Z}$ and $0\leq t\leq T_{2}$, we have
$2^{j}\|\dot{\Delta}_{j}\delta\eta_{n}(t)\|_{L^{2}}^{2}+\int_{0}^{t}2^{j}\|\nabla\dot{\Delta}_{j}\delta\eta_{n}\|_{L^{2}}^{2}d\tau\lesssim T_{2}\|\delta\tilde{H}_{n-1}\|_{C_{T_{2}}(\dot{B}^{-\frac{1}{2}}_{2,\infty})}^{2}.$
Thus, we obtain
$\|(\delta v_{n},\delta\eta_{n})\|_{C_{T_{2}}(\dot{B}^{\frac{1}{2}}_{2,\infty})}\lesssim T_{2}^{\frac{1}{2}}E_{0}\|(\delta v_{n},\delta\sigma_{n-1})\|_{C_{T_{2}}(\dot{B}^{\frac{1}{2}}_{2,\infty})}$
$\hskip 100.0pt+T_{2}^{\frac{1}{2}}\|(\delta\tilde{G}_{n-1},\delta\tilde{H}_{n-1})\|_{C_{T_{2}}(\dot{B}^{-\frac{1}{2}}_{2,\infty})}.$ (39)
The function $\delta\sigma_{n}$ satisfies the following initial value problem:
$\left\{\begin{aligned} &\partial_{t}\delta\sigma_{n}+v_{n-1}\cdot\nabla\delta\sigma_{n}=-\delta v_{n-1}\cdot\nabla\sigma_{n-1}+\delta\tilde{F}_{n-1}\ \ \mathrm{in}\ \ (0,T_{2})\times\mathbb{R}^{3},\\ &\delta\sigma_{n}|_{t=0}=0\ \ \ \mathrm{on}\ \ \ \mathbb{R}^{3},\end{aligned}\right$. (40)
where $\delta\tilde{F}_{n-1}=\tilde{F}_{n-1}-\tilde{F}_{n-2}$. By Lemma 2.3 and Lemma 2.4, we obtain
$2^{j}\|\dot{\Delta}_{j}\delta\sigma_{n}(t)\|_{L^{2}}^{2}+2^{j}\int_{0}^{t}\langle v_{n-1}\cdot\nabla\dot{\Delta}_{j}\delta\sigma_{n},\dot{\Delta}_{j}\delta\sigma_{n}\rangle d\tau$
$\lesssim\int_{0}^{t}2^{j}\|[\dot{\Delta}_{j},v_{n-1}\cdot\nabla]\delta\sigma_{n}\|_{L^{2}}^{2}$
$\hskip 50.0pt+\|-\delta v_{n-1}\cdot\nabla\sigma_{n-1}+\delta\tilde{F}_{n-1}\|_{\dot{B}^{\frac{1}{2}}_{2,\infty}}^{2}d\tau$
$\lesssim T_{2}E_{0}^{2}\|(\delta\sigma_{n},\delta v_{n-1})\|_{C_{T_{2}}(\dot{B}^{\frac{1}{2}}_{2,\infty})}^{2}+T_{2}\|\delta\tilde{F}_{n-1}\|_{C_{T_{2}}(\dot{B}^{\frac{1}{2}}_{2,\infty})}^{2}.$ (41)
By the estimate
$\langle v_{n-1}\cdot\nabla\dot{\Delta}_{j}\delta\sigma_{n},\dot{\Delta}_{j}\delta\sigma_{n}\rangle$ $=-\frac{1}{2}\langle\mathrm{div}v_{n-1},(\dot{\Delta}_{j}\delta\sigma_{n})^{2}\rangle$
$\lesssim\|\mathrm{div}v_{n-1}\|_{L^{\infty}}\|\dot{\Delta}_{j}\delta\sigma_{n}\|_{L^{2}}^{2},$
it follows from (3.2) that
$\|\delta\sigma_{n}\|_{C_{T_{2}}(\dot{B}^{\frac{1}{2}}_{2,\infty})}\lesssim T_{2}^{\frac{1}{2}}E_{0}\|(\delta\sigma_{n},\delta v_{n-1})\|_{C_{T_{2}}(\dot{B}^{\frac{1}{2}}_{2,\infty})}$
$\hskip 120.0pt+T_{2}^{\frac{1}{2}}\|\delta\tilde{F}_{n-1}\|_{C_{T_{2}}(\dot{B}^{\frac{1}{2}}_{2,\infty})}.$ (42)
By Lemma 2.4, we have
$\|\delta\tilde{F}_{n-1}\|_{C_{T_{2}}(\dot{B}^{\frac{1}{2}}_{2,\infty})}+\|(\delta\tilde{G}_{n-1},\delta\tilde{H}_{n-1})\|_{C_{T_{2}}(\dot{B}^{-\frac{1}{2}}_{2,\infty})}$
$\hskip 20.0pt\lesssim(E_{0}+E_{0}^{2})\|(\delta\sigma_{n-1},\delta v_{n-1},\delta\eta_{n-1})\|_{C_{T_{2}}(\dot{B}^{\frac{1}{2}}_{2,\infty})}.$ (43)
By (3.2), (3.2) and (3.2), if we take sufficiently small $T_{2}>0$, then the sequence $\{(\sigma_{n},v_{n},\eta_{n})\}_{n\geq 0}$ satisfies (37). Therefore, the sequence $\{(\sigma_{n},v_{n},\eta_{n})\}_{n\geq 0}$ is a Cauchy sequence in $C_{T_{2}}(\dot{B}^{1/2}_{2,\infty})$, is a bounded sequence in $C_{T_{2}}(\dot{H}^{1}\cap\dot{H}^{k})$ and satisfies (34), so that the convergence limit $(\rho,v,\theta)=(\rho_{\infty}+\sigma,v,\theta_{\infty}+\eta)$ of the sequence $\{(\rho_{n},v_{n},\eta_{n})\}$ is a solution of (1) which satisfies
$\|(\sigma,v,\eta)\|_{C_{T_{2}}(\dot{B}^{\frac{1}{2}}_{2,\infty})}\leq 2E_{0}.$
and
$\|\sigma\|_{C_{T_{2}}(L^{\infty})}\leq\tilde{c}_{2}\rho_{\infty}.$
3.3 Proof of Theorem 1.1
We now turn to prove Theorem 1.1. The following construction of time-periodic solutions is inspired by Valli [MR753158], which concerns a time-periodic problem for the isentropic compressible Navier-Stokes equation in a bounded domain.
Proof of Theorem 1.1.
By combining Proposition 3.1 and Proposition 3.9, there is small $\delta>0$ such that if initial data $(\rho_{0},v_{0},\theta_{0})=(\rho_{\infty},0,\theta_{\infty})+U_{0}$ and $f$ satisfy
$\|U_{0}\|_{\dot{B}^{\frac{1}{2}}_{2,\infty}\cap\dot{H}^{k}}+\|f\|_{C(\mathbb{R};\dot{B}^{\frac{1}{2}}_{2,\infty}\cap\dot{H}^{k-1})}\leq\delta,$
[…]
$\|U\|_{C(\dot{B}^{\frac{1}{2}}_{2,\infty}\cap\dot{H}^{k})}\lesssim\delta.$
Let $U^{*}$ be the global solution to (1) with zero initial data. Then, by taking $\delta>0$ sufficiently small, Proposition 3.1 and Proposition 3.2 imply that
$\|U^{*}\|_{C(\dot{B}^{\frac{1}{2}}_{2,\infty}\cap\dot{H}^{k})}\lesssim\delta$
and
$\|U^{*}(nT)-U^{*}(mT)\|_{\dot{B}^{1}_{2,\infty}\cap\dot{H}^{4}}\lesssim((n-m)T)^{-\frac{1}{4}}\delta$
for any $n,m\in\mathbb{Z}_{\geq 0}$ with $m\leq n$. Since the sequence $\{U^{*}(nT)\}_{n\in\mathbb{Z}_{\geq 0}}$ is bounded in $\dot{B}^{1/2}_{2,\infty}\cap\dot{H^{k}}$ and is a Cauchy sequence in $\dot{B}^{1}_{2,\infty}\cap\dot{H}^{4}$, there is $U^{*}_{\infty}\in\dot{B}^{1/2}_{2,\infty}\cap\dot{H}^{k}$ such that
$\lim_{n\to\infty}U^{*}(nT)=U^{*}_{\infty}\ \ \ \mathrm{in}\ \ \ \dot{B}^{1}_{2,\infty}\cap\dot{H}^{4}$
and
$\|U^{*}_{\infty}\|_{\dot{B}^{\frac{1}{2}}_{2,\infty}\cap\dot{H}^{k}}\lesssim\delta.$
Let $(\rho_{T},v_{T},\theta_{T})=(\rho_{\infty},0,\theta_{\infty})+U_{T}$ be a global solution of (1) with initial data $(\rho_{\infty},0,\theta_{\infty})+U^{*}_{\infty}$. Then, by Proposition 3.2, for any $n\in\mathbb{Z}_{>0}$, we have
$\|U_{T}(T)-U^{*}(nT))\|_{\dot{B}^{1}_{2,\infty}}\lesssim\|U^{*}_{\infty}-U^{*}((n-1)T)\|_{\dot{B}^{1}_{2,\infty}\cap\dot{H^{4}}}.$ (44)
Taking limit as $n\to\infty$ on both sides of (44) yields $U_{T}(T)=U^{*}_{\infty}$. By the local in time uniqueness in Proposition 3.9, we conclude that $(\rho_{T},v_{T},\theta_{T})$ is the desired time-periodic solution. The decay estimate (((ii))) follows immediately from Lemma 2.1 (iii), Lemma 2.2 and Proposition 3.2. ∎
Acknowledgment
I would like to thank Professor Yoshiyuki Kagei for his valuable comments. This work was supported by JSPS KAKENHI Grant Number JP23KJ0942.

### v2

@@
-Stability of time-periodic solutions to the Navier-Stokes-Fourier system
+Stability analysis of time-periodic solutions to the Navier-Stokes-Fourier system in 3D whole space
 Abstract
-We prove the existence and stability of a time-periodic solution to the Navier-Stokes-Fourier system in the three-dimensional whole space when a time-periodic external force is sufficiently small.
-The time decay estimate of the perturbation around the time-periodic solution is derived under the assumption that an initial perturbation is small and belongs to $L^{p}$ for some $1\leq p\leq 2$.
+This paper concerns the large-time behavior of perturbations around a time-periodic solution to the Navier-Stokes-Fourier system in the three-dimensional whole space.
+The time-periodic solution exists when a given external force is small enough.
+We derive the time-decay estimate of the perturbation under the assumption that an initial perturbation is sufficiently small.
+The time-space integral estimate for the linearized semigroup around the constant state in the Besov spaces is effectively applied in the proof.
 1 Introduction
 The motion of a compressible viscous and heat conducting fluid is described by the Navier-Stokes-Fourier system for the density $\rho$, the velocity $v=(v_{1},v_{2},v_{3})$ and the temperature $\theta$:
-$\left\{\begin{aligned} &\partial_{t}\rho+\mathrm{div}(\rho v)=0,\\ &\partial_{t}v+v\cdot\nabla v=\frac{\mu}{\rho}\Delta v+\frac{\mu+\mu^{\prime}}{\rho}\nabla\mathrm{div}\,v-\frac{\nabla P}{\rho}+f,\\ &\partial_{t}\theta+v\cdot\nabla\theta+\frac{\theta\partial_{\theta}P}{c_{V}\rho}\mathrm{div}\,v=\frac{\kappa}{c_{V}\rho}\Delta\theta+\frac{\Psi}{c_{V}\rho}.\end{aligned}\right$.
+$\left\{\begin{aligned} &\partial_{t}\rho+\mathrm{div}(\rho v)=0\ \ \ \mathrm{in}\ \ \ \mathbb{R}\times\mathbb{R}^{3},\\ &\partial_{t}v+v\cdot\nabla v=\frac{\mu}{\rho}\Delta v+\frac{\mu+\mu^{\prime}}{\rho}\nabla\mathrm{div}\,v-\frac{\nabla P}{\rho}+f\ \ \ \mathrm{in}\ \ \ \mathbb{R}\times\mathbb{R}^{3},\\ &\partial_{t}\theta+v\cdot\nabla\theta+\frac{\theta\partial_{\theta}P}{c_{V}\rho}\mathrm{div}\,v=\frac{\kappa}{c_{V}\rho}\Delta\theta+\frac{\Psi}{c_{V}\rho}\ \ \ \mathrm{in}\ \ \ \mathbb{R}\times\mathbb{R}^{3}.\end{aligned}\right$.
 (1)
-Here, $f$ is the external body force, $P$ is the pressure, $\mu$ and $\mu^{\prime}$ are the viscous coefficients, $\kappa$ is the heat conductivity coefficient, $c_{V}$ is the specific heat at constant volume and $\Psi=\Psi(v)$ is the dissipation function given by
+Here, $f=f(t,x)$ is the external body force, $P$ is the pressure, $\mu$ and $\mu^{\prime}$ are the viscous coefficients, $\kappa$ is the heat conductivity coefficient, $c_{V}$ is the specific heat at constant volume and $\Psi=\Psi(v)$ is the dissipation function given by
 $\Psi=\frac{\mu}{2}(\partial_{i}v^{j}+\partial_{j}v^{i})^{2}+\mu^{\prime}(\partial_{i}v^{i})^{2}.$
@@
 $\partial_{\rho}P>0\ \ \ \mathrm{and}\ \ \ \partial_{\theta}P>0.$
-The aim of this paper is twofold.
-The first is to investigate the existence of time-periodic solutions to the Navier-Stokes-Fourier system in $\mathbb{R}^{3}$ subject to a given small time-periodic external force.
-The second is to examine the stability of the time-periodic solutions with respect to external forces and initial perturbations, respectively.
 Let $f$ be a time-periodic force with time-period $T>0$, that is,
@@
 (2)
-We consider the time-periodic problem (2) around the constant state
+In this paper, we consider the time-periodic problem (2) around the constant state
 $(\rho_{\infty},0,\theta_{\infty})\ \ \ \mathrm{with}\ \ \ \rho_{\infty},\theta_{\infty}>0.$
-The time-periodic problem of the Navier-Stokes-Fourier system was considered by Ma, Ukai and Yang [MR2595722] in $\mathbb{R}^{N}$, $N\geq 5$.
-They proved the existence and stability result for the time-periodic problem by effectively applying the energy method in the inhomogeneous Sobolev space of $L^{2}$ type, but the dimensional assumption $N\geq 5$ is needed to control the $L^{2}$ norm of the time-periodic solution.
-Indeed, since the diffusion terms
-$\mu\Delta v+(\mu+\mu^{\prime})\nabla\mathrm{div}v\ \ \ \mathrm{and}\ \ \ \frac{\kappa}{c_{V}}\Delta\theta$
-are classified as elliptic operators, the spatial decay orders of the velocity $v_{T}$ and the temperature $\theta_{T}-\theta_{\infty}$ are expected at most $O(1/|x|^{N-2})$ as $|x|\to\infty$, respectively, by comparing the spatial decay order of the Newtonian potential.
-See also the asymptotic expansion result for the incompressible Navier-Stokes equation in [MR1794517].
-If the dimension $N=3$, then the expected spatial decay order is $O(1/|x|)$, which does not belong to $L^{2}$, so that this slow spatial decay suggests a need for a new function framework to construct a time-periodic solution in dimension $N=3$.
-In order to establish a function space setting that admits slow spatial decay, such as $1/|x|$, we introduce the homogeneous Besov spaces as follows.
-Let $s\in\mathbb{R}$ and $1\leq r\leq\infty$.
-Choose $\phi\in C^{\infty}$ supported in the annulus $\mathcal{C}=\{\xi\in\mathbb{R}^{3}\mid 3/4\leq|\xi|\leq 8/3\}$ such that
-$\sum_{j\in\mathbb{Z}}\phi(2^{-j}\xi)=1\ \ \ \mathrm{for}\ \ \ \xi\neq 0.$
-We define dyadic blocks $(\dot{\Delta}_{j})_{j\in\mathbb{Z}}$ by the Fourier multiplier
-$\dot{\Delta}_{j}u\equiv\mathcal{F}^{-1}[\phi(2^{-j}\cdot)\hat{u}$].
-Then, the homogeneous Besov space $\dot{B}^{s}_{2,r}=\dot{B}^{s}_{2,r}(\mathbb{R}^{3})$ is defined by
-$\dot{B}^{s}_{2,r}=\{u\in\mathcal{S}^{\prime}/\mathcal{P}\mid\|u\|_{\dot{B}^{s}_{2,r}}<\infty\}$
-with norm
-$\|u\|_{\dot{B}^{s}_{2,r}}:=\|\{2^{js}\|\dot{\Delta}_{j}u\|_{L^{2}}\}_{j\in\mathbb{Z}}\|_{\ell^{r}(\mathbb{Z})},$
-where $\mathcal{S}^{\prime}$ is the set of all Schwartz distributions and $\mathcal{P}$ is the polynomial ring over $\mathbb{R}$.
-We shall employ a function space to construct the time-periodic solution as the space $\dot{B}^{1/2}_{2,\infty}\cap\dot{H}^{k}$, $k\in\mathbb{Z}_{\geq 5}$, which spaces allow for slow decay, such as $1/|x|$, and have good compatibility with the energy method, since the space $\dot{B}^{1/2}_{2,\infty}\cap\dot{H}^{k}$ is $L^{2}$-based space.
-Here, we mention that the space $\dot{B}^{1/2}_{2,1}\cap\dot{H}^{3}$ was employed for the perturbation analysis of stationary solutions to the isentropic compressible Navier-Stokes equations in [MR4803477].
-Let us state our main theorem, which shows the existence and stability of the time-periodic problem (2).
+The time-periodic problem has been extensively studied for the incompressible Navier-Stokes equation; we refer here to Maremonti [12], Kozono-Nakao [9], Yamazaki [17], Galdi-Silvestre [5], Kobayashi [8] and Kučera [10] among many others.
+In particular, Yamazaki [17] derived time-decay estimates for perturbations around a uniformly bounded in time solution, including a time-periodic solution, to the incompressible Navier-Stokes equation in unbounded domains such as the whole space, the half-space, or exterior domains of dimension $n\geq 3$ under the assumption of smallness for both the external force and the initial perturbation.
+For the isentropic compressible Navier-Stokes equation, the existence and stability of the time-periodic solution were studied by Valli [16] in a 3D bounded domain, and by Kagei-Tsuda [6], Tsuda [15] in $\mathbb{R}^{n}$, $n\geq 3$.
+For the non-isentropic case, Ma, Ukai and Yang [11] proved the existence and stability result for the compressible Navier-Stokes system in $\mathbb{R}^{n}$, $n\geq 5$ and Tsuda proved the existence and stability result for the compressible Navier-Stokes-Korteweg system in $\mathbb{R}^{3}$.
+We also mention that Feireisl et al.
+[4] show the existence of at least one weak time periodic solution to the Navier-Stokes-Fourier system in a 3D bounded domain for large data.
+The time decay estimate of the perturbation around a time-periodic solution was derived by Ma, Ukai and Yang [11] for the Navier-Stokes-Fourier system under the smallness assumptions.
+They prove the decay estimate
+$\|(\rho-\rho_{T},v-v_{T},\theta-\theta_{T})(t)\|_{H^{N-1}}$
+$\hskip 50.0pt\lesssim(1+t)^{-\frac{n}{4}}\|(\rho_{0}-\rho_{T}(0),v_{0}-v_{T}(0),\theta_{0}-\theta_{T}(0))\|_{L^{1}\cap H^{N-1}},$ (3)
+where $N\in\mathbb{Z}_{\geq n+2}$ and $H^{N-1}$ is the inhomogeneous Sobolev space defined in the end of this section.
+Here, $(\rho,v,\theta)$ is a global solution to (1) in $\mathbb{R}^{n}$, $n\geq 5$ with initial data $(\rho_{0},v_{0},\theta_{0})$ and $(\rho_{T},v_{T},\theta_{T})$ is the corresponding time-periodic solution.
+The decay rate in (1) corresponds to that of the heat semigroup.
+However, the proof of decay estimate in [11] relies on the dimensional assumption $n\geq 5$.
+In this paper, we aim to establish the decay estimates for perturbations around a time-periodic solution to the Navier-Stokes-Fourier system in the physically important three-dimensional case.
+Let us state our main theorem, which derives the decay estimate for the perturbation.
 Theorem 1.1.
@@
 $\|f\|_{C(\mathbb{R};\dot{B}^{-\frac{3}{2}}_{2,\infty}\cap\dot{H}^{k-1})}\leq\delta,$
-then the following assertions hold.
+then the following assertions hold. Here, the space $\dot{B}_{2,\infty}^{\frac{1}{2}}$ is the Besov space defined in Section $2$ below.
   1. ((i))
@@
     for any $3/2+\epsilon\leq s\leq 3/2-\epsilon$ with s/2+3/2(1/p1/2)>0s/2+3/2(1/p-1/2)>0.
-The time-periodic problem for the incompressible Navier-Stokes equation has been extensively studied; we refer here to Maremonti [MR1107016], Kozono-Nakao [MR1373173], Yamazaki [MR1777114], Galdi-Silvestre [MR2603768], Kobayashi [MR2589952] and Kučera [MR2610568] among many others.
-In particular, Yamazaki [MR1777114] derived time-decay estimates for perturbations around a time-periodic solution to the incompressible Navier-Stokes equation in unbounded domains such as the whole space, the half-space, or exterior domains of dimension $N\geq 3$.
-For the isentropic compressible Navier-Stokes equation, the existence and stability of the time-periodic solution were studied by Valli [MR753158] in a 3D bounded domain, and by Kagei-Tsuda [MR3274764], Tsuda [MR3437858] in $\mathbb{R}^{N}$, $N\geq 3$.
-The existence and stability of the time-periodic solution to the Navier-Stokes-Fourier system was studied by Ma, Ukai and Yang [MR2595722] in $\mathbb{R}^{N}$, $N\geq 5$.
-We also mention that Feireisl et al.
-[MR2917121] show the existence of at least one weak time periodic solution of the Navier-Stokes-Fourier system in a 3D bounded domain for the specific pressure.
+We note that the decay rate obtained in (((ii))) coincides with that of the heat semigroup.
+The employment of the Besov space is motivated by the need to control the slow spatial decay of the time-periodic solution.
+Indeed, since the diffusion terms
+$\mu\Delta v+(\mu+\mu^{\prime})\nabla\mathrm{div}v\ \ \ \mathrm{and}\ \ \ \frac{\kappa}{c_{V}}\Delta\theta$
+are classified as elliptic operators, the spatial decay orders of the velocity $v_{T}$ and the temperature $\theta_{T}-\theta_{\infty}$ are expected at most $O(1/|x|)$ as $|x|\to\infty$, respectively, by comparing the spatial decay order of the Newtonian potential in $\mathbb{R}^{3}$.
+This slow decay suggests that the time-periodic solution does not belong to $L^{2}(\mathbb{R}^{3})$.
+Previously, Ma, Ukai and Yang [11] derived the decay estimate (1) in dimension $n\geq 5$ by combining the analysis using the linearized semigroup around the constant state $(\rho_{\infty},0,\theta_{\infty})$ and the energy estimates in $H^{N-1}$, $N\in\mathbb{Z}_{\geq n+2}$.
+In their work, the dimensional assumption $n\geq 5$ is used to control the $L^{2}$ norm of the time-periodic solution.
+To establish the decay estimate of the perturbation in dimension $n=3$, we develop the framework of perturbation analysis in the function spaces $\dot{B}^{1/2}_{2,\infty}\cap\dot{H}^{k}$, $k\geq\mathbb{Z}_{\geq 5}$, which spaces allow for slow decay, such as $1/|x|$.
+In addition, the spaces $\dot{B}^{1/2}_{2,\infty}\cap\dot{H}^{k}$, $k\geq\mathbb{Z}_{\geq 5}$ have good compatibility with the energy method developed by Matsumura and Nishida [14] for constructing strong solutions to the compressible Navier-Stokes-Fourier system, since the spaces are $L^{2}$-based space.
+Here, we mention that the space $\dot{B}^{1/2}_{2,1}\cap\dot{H}^{3}$ was used for the perturbation analysis of stationary solutions to the isentropic compressible Navier-Stokes equations in [3].
+The difficulty in proving Theorem 1.1 arises from the convection terms in the Navier-Stokes-Fourier system.
 Since the spatial decay order of the time-periodic solution is expected to be at most O(1/|x|){O}(1/|x|), the convection term $v\cdot\nabla v$ appearing in the Navier-Stokes equations is expected to exhibit the same spatial decay order, ${O}(1/|x|^{3})$, as the viscous term $\mu\Delta v$.
-Thus, we need to treat the convection terms carefully to drive the existence of a time-periodic solution even if the given external force is small enough.
-For this reason, Yamazaki and Tsuda transform the convection term $v\cdot\nabla v$ to the divergence forms $\mathrm{div}(v\otimes v)$ and $\mathrm{div}(\rho v\otimes v)$, respectively, by using the mass conservation low, and then establish a priori estimate for the solutions.
-In the Navier-Stokes-Fourier system case, the additional convection term $v\cdot\nabla\theta$ and the non-divergence form term $(\theta\partial_{\theta}P/c_{V}\rho)\mathrm{div}v$ appear in the equation of the temperature in (1).
+Thus, we need to treat the convection terms carefully to prove the existence of a time-periodic solution and to derive the decay estimate for its perturbations even if the given external force is small enough.
+For this reason, Yamazaki and Tsuda transform the convection term $v\cdot\nabla v$ to the divergence forms $\mathrm{div}(v\otimes v)$ and $\mathrm{div}(\rho v\otimes v)$, respectively, by using the mass conservation low, and then establish a priori bound for the solutions to the Navier-Stokes equations.
+However, in the Navier-Stokes-Fourier system case, the additional convection term $v\cdot\nabla\theta$ and the non-divergence form term $(\theta\partial_{\theta}P/c_{V}\rho)\mathrm{div}v$ appear in the equation of the temperature in (1).
 In Section 3 below, to establish a priori estimate in the homogeneous Besov spaces, we shall introduce the modified energy functional
@@
 Except for the external force terms and the higher-order terms, all nonlinear terms in (6) are written in divergence forms or potential forms.
-This transform is employed to establish a priori estimates for the initial value problem (1) in the proofs of Propositions 3.5 and 3.6 below.
-The proof of the time-decay estimate (((ii))) is inspired by the strategy employed in [MR4803477].
-The key step of the proof is to derive the time-weighted estimate for the Duhamel integral:
+This transform is used to establish a priori estimates for the initial value problem (1) in the proofs of Propositions 3.5 and 3.6 below.
+The proof of the time-decay estimate (((ii))) is inspired by the strategy used in [3].
+The proof will be carried out by writing the perturbation $U=(\sqrt{p_{1}}\sigma,\ m,\ \sqrt{c_{V}/\theta_{\infty}}\mathcal{E})$ as the Duhamel form:
+$U(t)=e^{tA}U(0)+\int_{0}^{t}e^{\tau A}\Phi(t-\tau)d\tau$
+where $t>0$,
+$A=\left[\begin{matrix}0&-\sqrt{p_{1}}\mathrm{div}&0\\ -\sqrt{p_{1}}\nabla&\frac{\mu}{\rho_{\infty}}\Delta+\frac{\mu+\mu^{\prime}}{\rho_{\infty}}\nabla\mathrm{div}&-\frac{p_{2}}{\rho_{\infty}}\sqrt{\frac{\theta_{\infty}}{c_{V}}}\nabla\\ 0&-\frac{p_{2}}{\rho_{\infty}}\sqrt{\frac{\theta_{\infty}}{c_{V}}}\mathrm{div}&\frac{\kappa}{c\rho_{\infty}}\Delta\end{matrix}\right],$
+and $\Phi$ denotes the remainder terms that contain the linear terms such as $\mathrm{div}(v_{T}\otimes m)$. The key step of the proof is to derive the time-weighted estimate for the Duhamel integral:
 $\sup_{0\leq t\leq\widetilde{T}_{1}}(1+t)^{\frac{s-s_{0}}{2}}\left\|\int_{0}^{t}e^{\tau A}\Phi(t-\tau)d\tau\right\|_{\dot{B}^{s}_{2,\infty}}$
-$\hskip 50.0pt\lesssim\sup_{s_{1}\leq s\leq\frac{3}{2}-\epsilon}\ \sup_{0\leq t\leq\widetilde{T}_{1}}(1+t)^{\frac{s-s_{0}}{2}}\|\Phi(t)\|_{\dot{B}^{s-2}_{2,\infty}\cap\dot{B}^{\frac{3}{2}}_{2,\infty}}$,
-where $\epsilon>0$, $3/2\leq s_{0}\leq 1/2$, $s_{1}=\max\{s_{0},0\}$, $\max\{s_{0},-3/2+\epsilon\}\leq s\leq 3/2-\epsilon$ and $e^{tA}$ is the linearized semigroup of (5) around the constant state which defined in (12) below.
-A similar analysis is found in Yamazaki [MR1777114].
+$\hskip 50.0pt\lesssim\sup_{s_{1}\leq\tilde{s}\leq\frac{3}{2}-\epsilon}\ \sup_{0\leq t\leq\widetilde{T}_{1}}(1+t)^{\frac{\tilde{s}-s_{0}}{2}}\|\Phi(t)\|_{\dot{B}^{\tilde{s}-2}_{2,\infty}\cap\dot{B}^{\frac{3}{2}}_{2,\infty}}$,
+where $\epsilon>0$, $3/2\leq s_{0}\leq 1/2$, $s_{1}=\max\{s_{0},0\}$, $\max\{s_{0},-3/2+\epsilon\}\leq s\leq 3/2-\epsilon$.
+By virtue of rewriting the convection term as the divergence form, the bilinear estimate in Lemma 2.5 below allow us to prove the estimates such as
+$\|\mathrm{div}(v_{T}\otimes m)\|_{\dot{B}^{\tilde{s}-2}_{2,\infty}\cap\dot{B}^{\frac{3}{2}}_{2,\infty}}\lesssim\|v_{T}\|_{\dot{B}^{\frac{1}{2}}_{2,\infty}\cap\dot{H}^{3}}\|m\|_{\dot{B}^{\tilde{s}}_{2,\infty}\cap\dot{H}^{3}},\quad s_{1}\leq\tilde{s}\leq 3/2-\epsilon,$
+so that if the time-periodic solution is sufficiently small and the high-frequency norm $\dot{H}^{3}$ of the perturbation is well controlled, then the linear operator $\mathrm{div}(v_{T}\otimes\cdot)$ can be treated as a simple perturbation of the operator $A$.
+The time-weighted estimate in the high-frequency norm is proved by using the energy method developed by Matsumura and Nishida [14] in Proposition 3.8 below.
 The outline of this paper is as follows: In Section 2, we present some lemmas related to the homogeneous Besov space, which will be frequently used in the subsequent sections.
@@
 (Cf.
-[MR1777114], [MR4803477].) The proof of the estimates in the homogeneous Sobolev norms is established by the energy method.
+[17], [3].) The proof of the estimates in the homogeneous Sobolev norms is established by the energy method.
 The proof of the local existence is carried out by constructing the approximating sequence of the solution.
 The approximating sequence is constructed as the solution of the transport equation for the density and as the solution of the quasilinear parabolic equation for the velocity and the temperature, since the density causes the quasilinear effect in the equations of the velocity and the temperature in (1).
-By virtue of employment in this construction, we obtain the local existence result without the smallness assumption for initial data, unlike the local existence result in [MR4803477].
+By virtue of employment in this construction, we obtain the local existence result without the smallness assumption for initial data, unlike the local existence result in [3].
 Finally, the proof of Theorem 1.1 is given in Subsection 3.3.
@@
 In particular, if $T_{0}=\infty$, we simply write $L^{p}(Z)$ and $C(Z)$ by omitting the subscript.
+For any $s\in\mathbb{R}$, $H^{s}=H^{s}(\mathbb{R}^{3})$ denotes the inhomogeneous Sobolev space by
+$H^{s}=\{u\in\mathcal{S}^{\prime}\mid(1+|\cdot|^{2})^{\frac{s}{2}}\hat{u}\in L^{2}\}$
+with norm
+$\|u\|_{H^{s}}=\|(1+|\cdot|^{2})^{\frac{s}{2}}\hat{u}\|_{L^{2}}.$
 2 Preliminary
-In this section, we present several lemmas related to the homogeneous Besov space, which are frequently used in this paper.
-The following basic properties for the homogeneous Besov space are proved in [MR2768550, Lemma 2.1, Proposition 2.22, Theorem 2.25 and Proposition 2.29].
+In this section, we present the definition and several lemmas for the homogeneous Besov space, which are frequently used in this paper.
+The following basic properties for the homogeneous Besov space are proved in [1, Lemma 2.1, Proposition 2.22, Theorem 2.25 and Proposition 2.29].
+Definition 2.1.
+Let $s\in\mathbb{R}$ and $1\leq r\leq\infty$. Choose $\phi\in C^{\infty}$ supported in the annulus $\mathcal{C}=\{\xi\in\mathbb{R}^{3}\mid 3/4\leq|\xi|\leq 8/3\}$ such that
+$\sum_{j\in\mathbb{Z}}\phi(2^{-j}\xi)=1\ \ \ \mathrm{for}\ \ \ \xi\neq 0.$
+We define dyadic blocks $(\dot{\Delta}_{j})_{j\in\mathbb{Z}}$ by the Fourier multiplier
+$\dot{\Delta}_{j}u\equiv\mathcal{F}^{-1}[\phi(2^{-j}\cdot)\hat{u}$].
+Then, the homogeneous Besov space $\dot{B}^{s}_{2,r}=\dot{B}^{s}_{2,r}(\mathbb{R}^{3})$ is defined by
+$\dot{B}^{s}_{2,r}=\{u\in\mathcal{S}^{\prime}/\mathcal{P}\mid\|u\|_{\dot{B}^{s}_{2,r}}<\infty\}$
+with norm
+$\|u\|_{\dot{B}^{s}_{2,r}}:=\|\{2^{js}\|\dot{\Delta}_{j}u\|_{L^{2}}\}_{j\in\mathbb{Z}}\|_{\ell^{r}(\mathbb{Z})},$
+where $\mathcal{P}$ is the polynomial ring over $\mathbb{R}$.
+The following lemma shows the basic properties of the Besov space.
 Lemma 2.2.
@@
     $\lim_{n\to\infty}u_{n}=u\ \ \mathrm{in}\ \ \mathcal{S}^{\prime}\ \ \ \mathrm{and}\ \ \ \|u\|_{\dot{B}^{s}_{2,r}\cap\dot{B}_{2,\tilde{r}}^{\tilde{s}}}\lesssim\liminf_{n\to\infty}\|u_{n}\|_{\dot{B}^{s}_{2,r}\cap\dot{B}_{2,\tilde{r}}^{\tilde{s}}}$.
-The following embedding property holds. (Cf. [MR2768550, Proposition 2.20, Propositon 2.39 and Theorem 2.40].)
+The following embedding property holds. (Cf. [1, Proposition 2.20, Propositon 2.39 and Theorem 2.40].)
 Lemma 2.3.
@@
 For any $2\leq q\leq\infty$, the space $\dot{B}^{3(1/2-1/q)}_{2,1}$ is continuously embedded in the space $L^{q}$.
-The following commutator estimate is the special case of [MR2768550, Lemma 2.100 and Remark 2.102].
+The following commutator estimate is the special case of [1, Lemma 2.100 and Remark 2.102].
 Lemma 2.4.
@@
 for any $1\leq k\leq 3$, where $u$ and $v$ are scalar functions.
-We state the bilinear estimate in the homogeneous Besov space. The following lemma is a direct consequence of [MR2768550, Theorem 2.47 and Theorem 2.52].
+We state the bilinear estimate in the homogeneous Besov space. The following lemma is a direct consequence of [1, Theorem 2.47 and Theorem 2.52].
 Lemma 2.5.
@@
 $\|uv\|_{\dot{B}_{2,\infty}^{s_{1}+s_{2}-\frac{3}{2}}}\lesssim\|u\|_{\dot{B}_{2,1}^{s_{1}}}\|v\|_{\dot{B}_{2,\infty}^{s_{2}}}.$
-We use the following lemmas regarding the composition of functions. (Cf. [MR2768550, Theorem 2.61].)
+We use the following lemmas regarding the composition of functions. (Cf. [1, Theorem 2.61].)
 Lemma 2.6.
@@
 3 Existence and stability of time-periodic solution
-This section is devoted to the proof of Theorem 1.1. In order to prove Theorem 1.1, we shall investigate the initial value problem of (1):
+This section is devoted to the proof of Theorem 1.1. We shall investigate the initial value problem of (1):
 $\left\{\begin{aligned} &\partial_{t}\rho+\mathrm{div}(\rho v)=0,\\ &\partial_{t}v+v\cdot\nabla v=\frac{\mu}{\rho}\Delta v+\frac{\mu+\mu^{\prime}}{\rho}\nabla\mathrm{div}\,v-\frac{\nabla P}{\rho}+f,\\ &\partial_{t}\theta+v\cdot\nabla\theta+\frac{\theta\partial_{\theta}P}{c_{V}\rho}\mathrm{div}\,v=\frac{\kappa}{c_{V}\rho}\Delta\theta+\frac{\Psi}{c_{V}\rho},\\ &(\rho,v,\theta)|_{t=0}=(\rho_{0},v_{0},\theta_{0})\end{aligned}\right$.
@@
 ∎
-The following lemma will be used to derive the time-weighted estimate. The proof of Lemma 3.4 below is the same as in [MR4803477, Proposition 5.2].
+The following lemma will be used to derive the time-weighted estimate. The proof of Lemma 3.4 below is the same as in [3, Proposition 5.2].
 Lemma 3.4.
@@
 We now turn to prove Theorem 1.1.
-The following construction of time-periodic solutions is inspired by Valli [MR753158], which concerns a time-periodic problem for the isentropic compressible Navier-Stokes equation in a bounded domain.
+The following construction of time-periodic solutions is inspired by Valli [16], which concerns a time-periodic problem for the isentropic compressible Navier-Stokes equation in a bounded domain.
 Proof of Theorem 1.1.
@@
 By the local in time uniqueness in Proposition 3.9, we conclude that $(\rho_{T},v_{T},\theta_{T})$ is the desired time-periodic solution.
-The decay estimate (((ii))) follows immediately from Lemma 2.1 (iii), Lemma 2.2 and Proposition 3.2.
+The decay estimate (((ii))) in Theorem 1.1 follows immediately from Lemma 2.2 (iii), Lemma 2.3 and Proposition 3.2.
 ∎